### v1

# System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity

Abstract
Generalized Nash equilibria are used in multi-agent control applications to model strategic interactions between agents that are coupled in the cost, dynamics, and constraints. We study the properties of open-loop GNE trajectories from a system-theoretic perspective. We show how strict dissipativity generates the turnpike phenomenon in GNE solutions. Moreover, we establish a converse turnpike result, i.e., the implication from turnpike to strict dissipativity. We derive conditions under which the steady-state GNE is the optimal operating point and, using a game value function, we give a local characterization of the geometry of storage functions. Finally, we design linear terminal penalties that ensure GNE open-loop trajectories converge to and remain at the steady-state GNE. These connections provide the foundation for future system-theoretic analysis of GNEs similar to those existing in optimal control.
keywords:
dynamic games, optimal control, Nash equilibrium problems, multi-agent systems, dissipativity, turnpike property, non-cooperative game theory
thanks: The authors are with the Automatic Control Lab, ETH Zürich, Physikstrasse 3, 8092 Zürich, Switzerland and with the Institute of Control Systems, Hamburg University of Technology, Harburger Schloßstraße 22a, 21079 Hamburg, Germany. This work was supported by the Swiss National Science Foundation under the NCCR Automation (grant 51NF40 225155).
1 Introduction
Control is at the core of many multi-agent applications managing interactions between strategic agents with conflicting objectives and coupled action spaces. Generalized Nash equilibria (GNEs) have emerged as a promising solution concept at the intersection of game theory and optimal control [Facchinei and Kanzow, 2009]. They naturally enable to define per-agent optimal control problems that are coupled, in a game-theoretic sense, in the objective, constraints, and dynamics. GNEs have been applied to resource allocation problems (e.g., energy [Atzeni et al., 2012, Hall et al., 2022], transportation [Bassanini et al., 2002], telecommunications [Pavel, 2012]) and competitive settings (e.g., autonomous driving [Dreves and Gerdts, 2017, Le Cleac’h et al., 2022], supply chains [Hall et al., 2024a]).
The prospect for applications has catalyzed extensive research on efficient GNE-seeking algorithms and their convergence guarantees [Facchinei et al., 2010, Belgioioso et al., 2022]. The resulting solution sets have been studied in detail [Kulkarni and Shanbhag, 2012, Nabetani et al., 2009], as well as their efficiency [Kulkarni, 2019] and fairness properties [Hall et al., 2025].
A fundamental step towards safe deployment of controllers is the characterization of their closed-loop properties. Yet, the behavior of finite-horizon trajectories resulting from GNEs remains poorly understood, even in the open-loop case. Previous works have applied dissipativity techniques to analyze the feedback mapping [Hall et al., 2024b] and drawn connections to dynamic variational inequalities [Benenati and Grammatico, 2024]. Despite this progress finite-horizon, GNE trajectories lack the rich system-theoretic characterizations developed for optimal control problems over the past decades, cf. [Willems, 1971, Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] and many others.
In classic optimal control, necessary first-order conditions are given by the maximum principle [Boltyanskii et al., 1960] while dynamic programming provides necessary and sufficient conditions [Bellman, 1954]. Moreover, the turnpike phenomenon is a similarity property of parametric optimal control problems with respect to varying initial conditions and horizon lengths. In essence, the phenomenon implies that for different initial conditions and horizon lengths the optimal trajectories approach the neighborhood of an optimal operating point (entry arc), remaining there (turnpike arc) and then diverge from it towards the end of the horizon (leaving arc). First observations date back to [Ramsey, 1928] and [von Neumann, 1938], while the name turnpike was coined by [Dorfman et al., 1958]. We refer to [McKenzie, 1976] and [Faulwasser and Grüne, 2022] for classic and modern literature overviews. A recent key development is the analysis of the links of the turnpike phenomenon with dissipativity properties [Grüne and Müller, 2016, Faulwasser et al., 2017], which is in turn linked to the dissipativity route to infinite-horizon optimal control which originated with [Willems, 1971], see [Faulwasser and Kellett, 2021] for recent nonlinear extensions. For results on turnpike analysis exploiting the structures of the maximum principle we refer to [Trélat and Zuazua, 2015]. In applications such as energy management, the turnpike effect or end-of-horizon effects lead, for instance, to myopic discharging of batteries [Hall et al., 2022, Vincent et al., 2020].
Interestingly, turnpike phenomena have also been observed in games in the Economics literature as early as the 1980s [Fershtman and Muller, 1986] and was later treated from a theoretical perspective for infinite-horizon open-loop games in continuous time [Carlson and Haurie, 1995] and discrete time [Carlson and Haurie, 1996] focusing on existence, uniqueness of equilibria and convergence to the turnpike. Extensions to differential games with coupled state constraints followed [Carlson and Haurie, 2000].
The turnpike property in games has recently received renewed attention. Turnpike properties for stochastic differential LQ games are studied in [Li et al., 2025] for the zero-sum setting and in [Cohen and Jian, 2025] for the non-zero sum setting. Multiple works derive turnpike properties for mean field games under a large population assumption [Cirant and Porretta, 2021, Carmona and Zeng, 2024, Ersland et al., 2025, Fedorov, 2025]. Further, turnpikes were observed in competitive supply chain games [Hall et al., 2024a].
This paper builds a bridge between the dissipativity route to turnpikes in optimal control and dynamic games. We establish the turnpike property for discrete-time finite-horizon GNE problems and study its implications on dissipativity of the GNE with respect to the steady-state GNE. Specifically, we make the following four contributions:
  1. 1.
    Demonstrating the crucial structural link between turnpike properties in parametric OCPs and parametric GNEs providing a foundation for system-theoretic analysis of GNEs.
  2. 2.
    Establishing that strict dissipativity implies the turnpike property, also in GNEs. Further, we derive a converse turnpike result, connecting the turnpike of the open-loop trajectory to dissipativity with respect to the steady-state GNE.
  3. 3.
    Providing an optimality-like interpretation of GNEs by considering a game value function and studying its local gradient structure. Additionally, we show that the storage function’s gradient equals the sum of agents’ dual multipliers at the steady-state GNE.
  4. 4.
    Designing mechanisms to suppress the leaving arc in GNE open-loop trajectories, using per-agent linear terminal penalties to ensure trajectories converge to and remain at the optimal operating point.
The remainder of this paper is organized as follows. Section 2 formulates the finite-horizon GNE problem for which we establish turnpike and dissipativity results in Section 3. We study connections between KKT systems of OCPs and GNEs in Section 4 and conclude the paper with simulation results in Section 5.
Notation
[…]
2 Problem Setting: Finite-Horizon Generalized Nash Equilibrium Problem
A group of self-interested agents $v\in\mathcal{V}:=\{1,\dots,M\}$ controls the following shared dynamics
$x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v})$ (1)
for a compact set of initial conditions $\mathbf{x}\in\mathbb{X}_{0}$. Each agent minimizes their objective function over a prediction horizon $N$ with states and inputs coupled to other agents. This constitutes a finite-horizon Generalized Nash Equilibrium problem (GNEP) as follows
$\displaystyle\min_{u^{v},\,x}$ $\;\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})$ (2a)
s.t. $x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v})\hskip 12.50002ptk\in\mathbb{Z}_{N}$ (2b)
$g(x_{k},u_{k}^{v},u^{-v}_{k})\leq 0,\hskip 25.00003ptk\in\mathbb{Z}_{N}$ (2c)
$G^{v}\,u_{k}^{v}\leq h^{v},\hskip 54.00002ptk\in\mathbb{Z}_{N}$ (2d)
$\;x_{0}=\mathbf{x}.$ (2e)
where $\mathbf{x}$ is the initial condition of system (1). We define the following per-agent and global action sets
[…]
and the cumulative cost
$J_{N}^{v}(x,u):=\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}).$
Decisions that jointly solve (2) are generalized Nash equilibria (GNE) [Facchinei and Pang, 2009, §2], as defined next.
Definition 1 (Generalized Nash equilibrium).
A joint decision $(x^{*},u^{*})$ is a GNE of (2) if $\forall v\in\mathcal{V}$
$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})$
$\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$. The corresponding solution set for fixed $N\in\mathbb{N}$ and $\mathbf{x}\in\mathbb{X}_{0}$ is denoted as
$(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\subset\mathbb{R}^{(N+1)n_{x}+Nn_{u}}.$ (5)
Intuitively, a GNE is a joint decision for which no agent $v\in\mathcal{V}$ can reduce its cost by unilaterally changing its own decision.
For the sake of concise exposition and in analogy to optimal control concepts, we refer to $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ as a GNE pair or an optimal game pair. Moreover, with slight abuse of notation we write
$(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$
to express that there exist $k\in\mathbb{Z}_{N}$ and $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ such that
$(x_{k},u_{k})=(x_{k}^{*},u_{k}^{*})\quad\text{with}\quad(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}).$
We want to study the convergence of the finite-horizon GNE trajectory with respect to a steady-state point. Thus, we introduce the game-theoretic equilibrium of the steady-state game next.
Definition 2 (Steady-state GNE).
The pair $(x_{s},u_{s})$ is called a steady-state GNE if it solves the steady-state GNEP
[…]
with the corresponding solution set $\mathcal{S}^{\text{\tiny GNE}}_{s}\subset\mathbb{R}^{n_{x}+n_{u}}.$
We analyze properties of the game pair of the GNE in (2) provided a solution exists. Thus, in the following we assume that $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $\mathcal{S}^{\text{\tiny GNE}}_{s}$ are nonempty $\forall\mathbf{x}\in\mathbb{X}_{0}$.
Remark 1 (Existence & computation).
The theoretical results which follow hold for the nonlinear GNE problem in (2). Existence of a GNE has been established under a variety of assumptions as summarized in [Dutang, 2013]. Efficient algorithms to compute solutions of nonlinear GNEPs are a topic of ongoing research, one example is given in [Le Cleac’h et al., 2022].
3 Dissipativity and Turnpike Relations for GNEPs
Turnpike and dissipativity properties are closely linked, providing system-theoretic insights for optimal control. In the following, we demonstrate that this link also exists for the GNE problem in (2). We first introduce a performance measure for the population of agents as follows
$\ell(x_{k},u_{k}):=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})$ (11)
and
$J_{N}(x,u):=\sum_{k=0}^{N-1}\ell(x_{k},u_{k})=\sum_{k=0}^{N-1}\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$
While using the sum of costs may seem unusual from a game-theoretic perspective, [Benenati and Grammatico, 2024] show that it can serve as a Lyapunov function for LQ games with shared constraints under suitable controllability assumptions. This motivates our choice in (11) and simplifies the subsequent theoretical analysis.
3.1 Dissipativity of GNEPs
Based on the joint performance measure, we can make a strict dissipativity assumption analogous to the ones existing in optimal control.
Definition 3 (Strict dissipativity of GNEPs).
Given a steady-state GNE $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$, the GNEP (2) is called strictly dissipative if for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ there exists a storage function $\Lambda:\mathbb{X}_{0}\to\mathbb{R}$ bounded from below and $\alpha_{\ell}\in\mathcal{K}$ such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$
$\Lambda(f(x_{k},u_{k}))-\Lambda(x_{k})\leq\\ -\alpha_{\ell}\left(\left\|\begin{matrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{matrix}\right\|\right)+s(x_{k},u_{k})$ (11)
holds for each $(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
Figure 1: Schematic of a state trajectory exhibiting the turnpike property for different horizon lengths.
[Willems, 1972] introduced equivalent dissipativity characterizations via the required supply and the available storage. In discrete time the available storage characterization can be traced back to [Byrnes and Lin, 1994]. For our further investigations, the latter one is very helpful. In terms of GNEPs, this well-known characterization reads as follows.
Theorem 1 (Available storage of GNEP ).
The GNEP (2) is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ and $\alpha_{\ell}\in\mathcal{K}_{\infty}$ if and only if
$\Lambda_{a}(\mathbf{x})=$ $\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x^{*},u^{*})\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}\sum_{k=0}^{N-1}-s(x_{k}^{*},u_{k}^{*})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)<\infty.$
Note that Definition 3 and Theorem 1 consider strict dissipativity only along game pairs resulting from the GNEP in (2). In optimal control, strict dissipativity is often assumed to hold along all feasible trajectories, see for instance [Grüne and Müller, 2016, Defn. 2.1], while [Faulwasser et al., 2017] also restricted the crucial inequality to hold only along optimal trajectories. However, in a game-theoretic setting we need to restrict the definition as the following shows.
Remark 2 (Why restrict dissipativity to $\mathcal{S}^{\text{\tiny GNE}}_{N}?$).
Denote the optimal steady state associated with (11) as
$(x^{\diamond},u^{\diamond})=$ $\operatorname*{arg\,min}_{\bar{x},\bar{u}}\;\ell(\bar{x},\bar{u})\;$ (12)
$\text{s.t.}\;\bar{x}=f(\bar{x},\bar{u}),\;g(\bar{x},\bar{u})\leq 0,\;G\,\bar{u}\leq h.$
Substituting it into (11) yields
$\ell(x_{s},u_{s})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{\diamond}-x_{s}\\ u^{\diamond}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x^{\diamond},u^{\diamond}),$
where $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$. This inequality, however, holds only if $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$ which implies that the steady-state GNE (10) has to coincide with the minimizer of (12), and thus be a solution to an optimization problem. In turn, this would require the game to be potential [Monderer and Shapley, 1996] and the potential function to be the sum of agents’ costs [Slade, 1994, Prp. 1]. More specifically, we see from the first-order optimality conditions that it would require
$\frac{\ell^{v}(x_{k},u_{k})}{\partial x_{k}}=\frac{\ell(x_{k},u_{k})}{\partial x_{k}}$ (13a)
$\frac{\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}=\frac{\ell(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}$ (13b)
and as $\ell(x_{k},u_{k})=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k})$ this would impose that all agents’ cost functions are the same up to a constant which does not hold for the GNEPs we consider in (2). It is a well known phenomenon that selfish agent behavior in games leads to solutions that are suboptimal compared to centrally minimizing the sum of costs (Price of Anarchy). [Kulkarni, 2019] analyze these efficiency losses for GNEs. Thus, in Definition 3 we require the dissipation inequality to hold only for trajectories resulting from the GNEP in (2).
The next result shows that strict dissipativity also provides an avenue to analyze which $(x_{S},u_{s})\in\mathcal{S}^{GNE}_{S}$ can appear in (11).
Lemma 2.
Let the strict dissipation inequality (11) hold for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$. Then
$\sum_{v\in\mathcal{V}}\ell^{v}(x_{s},u_{s}^{v})=:\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u}):=\sum_{v\in\mathcal{V}}\ell^{(}\bar{x},\bar{u}^{v})$
holds for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$.
Proof 3.1.
For contradiction, suppose there exists a better steady state solution $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$ with $\ell(x_{s},u_{s})>\ell(\bar{x},\bar{u})$. At steady state, the strict inequality (11) gives
$0<\alpha_{\ell}\left(\left\|\begin{matrix}\bar{x}-x_{s}\\ \bar{u}-u_{s}\end{matrix}\right\|\right)\leq\ell(\bar{x},\bar{u})-\ell(x_{s},u_{s}).$
Since $\alpha_{\ell}\in\mathcal{K}$, this allows writing $\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u})$ and thus shows the contradiction.
The last result requires (11) to hold on (10). This is needed, as without further analysis it is not clear that for $x_{0}=x_{s}$ the GNE control action at time $k=0$ is such that $u^{*}_{0}=u_{s}$. Indeed, the control action $u_{0}^{*}$ depends on the specific design of (2) and the specific horizon $N$. Corollary 10 in Section 4 presents a result that ensures $u^{*}_{0}=u_{s}$ if $x_{0}=x_{s}$.
We now formally introduce the considered turnpike property for our GNE problem in (2) with respect to the steady-state GNE.
Definition 4 (Measure turnpike in GNEPs).
The GNEP (2) exhibits the (measure) turnpike property at $(x_{s},u_{s})$ if for each $\varepsilon>0$ there exists $C>0$ such that $\forall N\in\mathbb{N},\mathbf{x}\in\mathbb{X}_{0}$ and for all game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ it holds that
$Q_{\varepsilon}:=\#\{k\in\mathbb{Z}_{N}|\,\|x_{k}-x_{s}\|\leq\varepsilon\}\geq N-\frac{C}{\alpha_{\ell}(\varepsilon)}.$ (14)
We next show that strict dissipativity of the GNEP with respect to $(x_{s},u_{s})$ implies the turnpike property for the input and state trajectories resulting from (2). The result is inspired by [Grüne, 2013, Thm. 5.3] and its proof.
Theorem 3 (Str. dissipativity $\Rightarrow$ turnpike).
If, for every $x\in\mathbb{X}_{0}$, the GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ in the sense of Definition 3, then all GNEP solutions with $J_{N}(x,u)\leq N\ell(x_{s},u_{s})+\delta$ exhibit the turnpike property at $(x_{s},u_{s})$.
Figure 2: Overview of implications between strict dissipativity, turnpike and optimal operation at the steady-state GNE.
Proof 3.2.
We define $C:=2\sup_{\mathbf{x}\in\mathbb{X}_{0}}|\Lambda(\mathbf{x})|<\infty$ and further we have that
$\tilde{J}_{N}(x,u)$ $=J_{N}(x,u)+\Lambda(\mathbf{x})-\Lambda(x_{N})$ (15a)
$\leq J_{N}(x,u)+C,$ (15b)
$\leq N\,\ell(x_{s},u_{s})+(\delta+C)$ (15c)
Also due to strict dissipativity we have that $\forall(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}),\,\forall\mathbf{x}\in\mathbb{X}_{0}$:
$\ell(x_{k},u_{k})+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))\geq$
$\ell(x_{s},u_{s})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\$(16)
For contradiction suppose that the trajectory does not fulfill the turnpike property and thus, $Q_{\varepsilon}<N-\frac{\delta+C}{\alpha_{\ell}{(\varepsilon})}$. This implies that there exists a number of $N-Q_{\varepsilon}>\frac{\delta+C}{\alpha_{\ell}(\varepsilon)}$ times steps in the set $\mathcal{N}\subseteq\{0,\dots,N-1\}$ for which the trajectory is more than $\varepsilon$ distance away from the steady state $x_{s}$, i.e., $\|x_{k}-x_{s}\|>\varepsilon,\;\forall k\in\mathcal{N}$.
$\tilde{J}_{N}(x,u)$ $=\sum_{k=0}^{N-1}[\ell(x_{k},u_{k}))+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))]$
[…]
$=N\ell(x_{s},u_{s})+Q_{\varepsilon}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)]$
$+\underbrace{(N-Q_{\varepsilon})\,\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)}_{\geq(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)}$
$\geq N\ell(x_{s},u_{s})+(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)$
$>N\ell(x_{s},u_{s})+(\delta+C)$
which clearly contradicts (15), thus the assertion.
Theorem 3 demonstrates that strict dissipativity of the GNEP (2) is not merely a theoretical property but gives direct insights into the behavior of the trajectory with respect to the steady-state GNE.
Remark 3 ( Turnpike with strict $x$ dissipativity).
Observe that—similar to its optimal control counterpart—the proof of Theorem 3 does not rely on $(x,u)$-dissipativity but only requires $x$-dissipativity with $\alpha_{\ell}(\|x-x_{s}\|)$. However, having strictness in $x$ and $u$ helps to derive converse turnpike results.
3.2 A converse turnpike result for GNEPs
The question arises if the converse also holds, i.e., whether the turnpike behavior in the open-loop trajectories implies strict dissipativity with respect to $(x_{s},u_{s})$. We recall the following “off-equilibrium” concept.
Definition 5 ($\varepsilon$-GNE [Chen et al., 2021]).
For a given $\varepsilon\geq 0$, $u^{*}$ is said to be an $\varepsilon$-generalized Nash equilibrium ($\varepsilon$-GNE) of (2) if $\forall v\in\mathcal{V}$ and $\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$ it holds:
$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})+\varepsilon,\;$
The solution set is denoted as $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})$. Particularly, when $\varepsilon=0$, $u^{*}$ is a GNE and $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})=\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
We make the following assumption which ensures that $(x_{s},u_{s})$ is the unique GNE in the set of $\rho$-GNEs centred at $x_{s}$ , $\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})$, which minimizes the performance measure $J(x,u)$ defined for the population of agents.
Assumption 1 (Local cost minimizer $(x_{s},u_{s})$).
There exists a constant $\rho$ and $\alpha_{\rho}\in\mathcal{K}$ such that $\forall(x,u)\in\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})\cap\mathcal{Z}(x_{s})$ and $\forall k\in\mathbb{N}$ it holds:
$\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x_{k},u_{k})-\ell(x_{s},u_{s}).$ (17)
Assumption 1 can be verified by employing the game-theoretic counterpart of Fiacco-like sensitivity analysis. Specifically, in the GNE setting formulated as quasi-variational inequalities, local uniqueness results exist under different sets of assumptions, see [Alphonse et al., 2021, Ch. 3] and [Noor et al., 2025, Lem. 7]. For variational GNEs, sensitivity analysis of variational inequalities can be exploited [Facchinei and Pang, 2009, Prop. 12.14]. The following result confirms a converse relation between turnpike and strict dissipativity.
Theorem 4 (Turnpike $\Rightarrow$ str. dissipativity).
Suppose at $(x_{s},u_{s})$ Assumption 1 and the turnpike property as per Definition 4 holds for the GNEP (2), then the GNEP is strictly dissipative for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ with respect to $(x_{s},u_{s})$.
Proof 3.3.
Consider $\alpha_{\rho}\in\mathcal{K}$ from Assumption 1 and note that for any $\rho$ there exist an $N\in\mathbb{N}$ such that $Q_{\varepsilon}>0$ as defined in (14). Consequently, we choose $\rho\geq\varepsilon$ and thus whenever we are in $Q_{\varepsilon}$ we can exploit Assumption 1.
Recall the available storage from Theorem 1
$\tilde{\Lambda}_{a}(\mathbf{x})=\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$ (18)
evaluated along the game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. Temporarily, consider some $N\in\mathbb{N}$ and a feasible initial condition $\mathbf{x}$. We split the time horizon into $Q_{\varepsilon}(N,\mathbf{x})$, the trajectory points close to $(x_{s},u_{s})$ and $Q_{out}(N,\mathbf{x})=\{0,\dots,N-1\}\backslash Q_{\varepsilon}(N,\mathbf{x})$ points outside of ball of radius $\varepsilon$. Evaluating (18) on $Q_{out}(N,\mathbf{x})$ gives
$\sum_{k\in Q_{out}(N,\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$
[…]
where $\hat{\alpha}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ and similarly $\hat{s}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}s(x_{k},u_{k})$. The Turnpike property implies that any trajectory, independent of $N$ and $\mathbf{x}$, only spends a finite amount of time outside of ball the of radius $\varepsilon$ of $x_{s}$ which allows us to establish the following bound:
[…]
Next, we evaluate the second part of the trajectory for $Q_{\varepsilon}(N,\mathbf{x})$ we have that $\|x_{k}-x_{s}\|\leq\rho,\;\forall k\in Q_{\varepsilon}(N,\mathbf{x})$ and given Assumption 1 it holds that:
$\sum_{Q_{\varepsilon}(N,\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})\leq 0$ (21)
Combining (3.3) and (21) we get that
[…]
3.3 Optimal operation at steady-state GNE
We have established in the previous section under which conditions the open-loop trajectory will spend most of its time at the turnpike, the steady-state GNE. However, this raises the question if the pair $(x_{s},u_{s})$ is also optimal in terms of the entire population performance measure $J(x,u)$. To investigate this, we define optimal steady state operation and suboptimal operation, following [Müller et al., 2015, Def. 2].
Definition 6 (Optimal operation of RHG).
The GNEP (2) is optimally operated at steady-state, if for each $\mathbf{x}\in\mathbb{X}_{0}$ and each game pair $(x,u)\in\mathcal{S}^{\text{\tiny NE}}_{N}(\mathbf{x})$ the following holds $\forall k\in\mathbb{I}_{\geq 0}$:
$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}\geq\ell(x_{s},u_{s}).$ (22)
it is suboptimally operated off steady state if (22) holds in a strict sense.
Optimal steady-state GNE operation implies that no feasible game pair of (2) can have a better (asymptotic) average performance than the performance of the steady-state GNE $(x_{s},u_{s})$, and suboptimal operation off steady-state means the performance is strictly worse or “passes by” $x_{s}$ infinitely often. Next, we relate strict dissipativity of (2) to optimal operation at $(x_{s},u_{s})$.
Proposition 1.
Suppose the set $\mathcal{S}_{\infty}^{GNE}(\mathbf{x})\not$$\emptyset$, then if GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ following Definition 3, it is optimally operated at the steady-state GNE $(x_{s},u_{s})$ and suboptimally operated off steady state.
Proof 3.4.
The proof follows along the lines of [Prop. 6.4][Angeli et al., 2012]. First, we note that $\sup_{x\in\mathbb{X}_{0}}|\Lambda(x)|<\infty$ which allows to state the following equality
$0$ $=\lim_{N\to+\infty}\frac{\Lambda(x_{N})-\Lambda(x_{0})}{N}$
$=\lim_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\Lambda(x_{k+1})-\Lambda(x_{k})}{N}$
$\leq\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)}{N}-\ell(x_{s},u_{s})$
From which it follows that
$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}$
$\geq\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\alpha_{\ell}\left(\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\|\right)}{N}+\ell(x_{s},u_{s}).$
We consider an infinite horizon feasible game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{\infty}(\mathbf{x})$ and distinguish two cases:
(i) If $x_{k}\not\to x_{s}$ as $N\to\infty$, then
$\liminf_{N\to+\infty}\frac{1}{N}\left(\sum_{k=0}^{N-1}\ell(x_{k},u_{k})\right)>\ell(x_{s},u_{s})$
due to the positive definiteness $\alpha_{\ell}(\cdot)$ proving suboptimal operation off steady state;
(ii) Else if $\liminf_{N\to+\infty}\|x_{k}-x_{s}\|=0$ and $\liminf_{N\to+\infty}\|u_{k}-u_{s}\|=0$ then
$\liminf_{N\to+\infty}(\ell(x_{k},u_{k}))/N)=\ell(x_{s},u_{s})$
proving optimal operation at steady state.
4 Optimality Interpretation of GNEs
In this section we aim to draw a connection between the GNE solution of (2) and the KKT conditions of optimal control problems. In order to do so, we state the following assumptions on the functions in the GNEP (2).
Assumption 2 (Differentiable problem data).
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ and constraint function $g(\cdot,\cdot,u^{-v})$ and the dynamics $f(\cdot,\cdot,u^{-v})$ are continuously differentiable in $x$ and $u^{v}$.
Assumption 3 (Convexity).
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ is convex and the set $\mathcal{Z}^{v}(\cdot,u^{-v})$ is closed and convex.
The convexity assumption is required to ensure that points which solve the individual agent’s KKT system are not merely stationary points (local minima, saddle points). KKT conditions of GNEs and relevant constraint qualifications have been studied in detail in [Bueno et al., 2019].
At each time step $k=0,\ldots,N-1$ consider
$L^{v}_{k}=\begin{bmatrix}1\\ \lambda^{v}_{k+1}\\ \mu_{k}^{v}\\ \nu_{k}^{v}\end{bmatrix}^{\top}\begin{bmatrix}\ell^{v}(x_{k},u_{k},u^{-v}_{k})\\ f(x_{k},u_{k}^{v},u_{k}^{-v})-x_{k+1}\\ g(x_{k},u_{k}^{v},u^{-v}_{k})\\ G^{v}u_{k}^{v}-h^{v}\end{bmatrix}$.
The Lagrangian of each agent associated with (2) can then be written as
$\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})=\sum_{k=0}^{N-1}L_{k}^{v}+(\lambda_{0}^{v})^{\top}[\mathbf{x}-x_{0}]$ (23)
This yields the following per-agent KKT system which needs to hold $\forall v\in\mathcal{V}$ and $k\in\mathbb{N}$
$x_{k+1}$ $=f(x_{k},u_{k}^{v},u_{k}^{-v})$ (24a)
$\lambda^{v}_{k}$ $=\ell_{x}^{v}+g_{x}^{\top}\mu^{v}_{k}+f_{x}^{\top}\lambda_{k+1}^{v},\quad$ (24b)
$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\mu^{v}_{k}+f_{u^{v}}^{\top}\lambda_{k+1}^{v}+(G^{v})^{\top}\eta_{k}^{v},$ (24c)
subject to the initial and boundary conditions
$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda^{v}_{N}=g_{x}^{\top}\mu^{v}_{N}$ (24d)
and the usual conditions of primal feasibility and complementarity slackness
$0\leq\mu_{k}^{v}$ $\quad\perp\quad-g^{v}(x_{k},u_{k}^{v},u_{k}^{-v})\geq 0,$ (24e)
$0\leq\eta_{k}^{v}$ $\quad\perp\quad h^{v}-G^{v}u^{v}_{k}\geq 0.$ (24f)
A well-established result from the GNE literature connects this KKT system to GNEs. The following result appeared as Thm. 4.6 in [Facchinei and Kanzow, 2009].
Theorem 5 (KKT conditions for GNEPs).
Given Assumption 2 then the following statements hold for the GNEP (2):
  1. (a)
    Let $(x^{*},u^{*})$ be an equilibrium of the GNEP at which all agents’ subproblems satisfy Slater’s constraint qualification. Then, $\forall v\in\mathcal{V}:\exists\,(\lambda^{v*},\mu^{v*},\eta^{v*})$ that together with $(x^{*},u^{*})$ solve the KKT system (24).
  2. (b)
    If $(x^{*},u^{*})$ and $(\lambda^{v*},\mu^{v*},\eta^{v*})_{v\in\mathcal{V}}$ solve the KKT system (24) and Ass. 3 holds, then $(x^{*},u^{*})$ is an equilibrium point of the GNEP.
Similar to [Faulwasser and Zanon, 2018], we study when a primal-dual solution to the steady-state GNEP (10) is also a solution for the dynamic counterpart (2). To this end, we consider the Lagrangian for each agent of the steady-state GNE problem in (10) given by
$\mathcal{L}_{s}^{v}(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})=\ell^{v}(\bar{x},\bar{u}^{v},\bar{u}^{-v})+(\bar{\mu}^{v})^{\top}\,g(\bar{x},\bar{u}^{v},\bar{u}^{-v})$
$+(\bar{\lambda}^{v})^{\top}[f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}]+(\bar{\eta}^{v})^{\top}(G^{v}\bar{u}^{v}-h^{v}).$
Computing the usual stationarity condition with respect to $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ gives the following set of conditions
$0$ $=f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x},$ (25a)
$\bar{\lambda}^{v}$ $=\ell_{x}^{v}+g_{x}^{\top}\bar{\mu}^{v}+f_{x}^{\top}\bar{\lambda}^{v},$ (25b)
$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\bar{\mu}^{v}+f_{u^{v}}^{\top}\bar{\lambda}^{v}+(G^{v})^{\top}\bar{\eta}^{v},$ (25c)
which together with
$0\leq\bar{\mu}^{v}$ $\quad\perp\quad-g^{v}(\bar{x},\bar{u}^{v},\bar{u}^{-v})\geq 0,$ (25d)
$0\leq\bar{\eta}^{v}$ $\quad\perp\quad h^{v}-G^{v}\bar{u}^{v}\geq 0,$ (25e)
form the KKT conditions for (10).
4.1 The game value function of GNEPs
In the analysis of optimal control problems, the value function plays a fundamental role. It allows to conclusively characterize optimality in Dynamic Programming, it can also be linked to the Maximum Principle. Specifically, in first-order optimality conditions of the Maximum Principle or its simpler Euler-Lagrange variants the gradient of the value function is related to the co-state/adjoint, i.e., to the dual variable linked to the equality constraints imposed by the dynamics. Given that (24) are from a per-agent perspective just first-order (KKT) conditions—though all agents consider the same dynamics and they have to agree on the same sate trajectory—one may wonder what is a GNEP-specific counterpart of the optimal control value function?
To this end, recall the construction of $\ell$ as the sum over all agent-specific costs $\ell^{v}$, cf. (11). Summing $\ell$ from $k=0$ to k=N1k=N-1 leads to the game value function which measures the performance of the agent population at $\mathbf{x}$
$\hskip-4.2679ptV_{N}(\mathbf{x}):=\sum_{k=0}^{N-1}\ell(x_{k}^{*},u^{*}_{k})=\sum_{v\in\mathcal{V}}\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$ (26)
Let $\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}$ denote an open ball of radius $\varepsilon$, centered at $\mathbf{x}$.
Theorem 6 (Sensitivity characterization).
Consider the GNEP (2), where all inequality constraints are neglected, and let Assumptions 23 hold. Suppose that for some $\varepsilon>0$ and for any perturbed initial condition
$x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$
we have $\mathcal{S}^{GNE}_{N}(\mathbf{x}+\xi)\not$$\emptyset$, i.e., the GNEP (2) admits a solution such that $V_{N}(\mathbf{x})$ is locally differentiable at $\mathbf{x}$.
Then
$\nabla V_{N}(\mathbf{x})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0},$ (27)
where $\lambda^{v}_{0}$ corresponds to the KKT solution tuple $(x,u,\lambda^{v},0,0)$ with $(x,u)\in\mathcal{S}^{GNE}_{N}(\mathbf{x})$.
Proof 4.1.
Consider the GNEP (24) without inequality constraints and subject to the perturbed initial condition
$x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}.$
For all $v\in\mathcal{V}$, let $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ denote the solution to the KKT conditions (24) with perturbed initial condition, which for all $\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$ are optimal due to Theorem 5.
Clearly, it holds that the solution for the unperturbed initial condition is given by
$(x^{*},u^{v,*},\lambda^{v})=(x(0),u^{v}(0),\lambda^{v}(0)).$
Computing the sensitivity of the Lagrangian (23) with respect to $\xi$ at $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ gives
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}\left(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)\right)=$ $\phantom{+}\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial x_{k}}\dfrac{\partial x_{k}}{\partial\xi}$
$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial u^{v}_{k}}\dfrac{\partial u^{v}_{k}}{\partial\xi}$
$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial\lambda^{v}_{k}}\dfrac{\partial\lambda^{v}_{k}}{\partial\xi}$
$+\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi},$
where $\frac{\partial x_{k}}{\partial\xi}$ is the sensitivity of state $x_{k}(\xi)$ to the perturbation of the initial condition $x_{0}=\mathbf{x}+\xi$. Likewise $\frac{\partial u^{v}_{k}}{\partial\xi}$ and $\frac{\partial\lambda^{v}_{k}}{\partial\xi}$ are sensitivities with respect to $\xi$.
Notice that in the KKT conditions (24) the optimality-like system (24a)–(24c) follows from
$\nabla_{x_{k}}L^{v}_{k}=0,\nabla_{u^{v}_{k}}L^{v}_{k}=0,\quad\text{and}\quad\nabla_{\lambda^{v}_{k}}L^{v}_{k}=0.$
Hence, for all $v\in\mathcal{V}$, we obtain
$\left.\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))\right|_{\xi=0}$ $=\left.\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi}\right|_{\xi=0}$
$=\lambda^{v}(0)^{\top}.$ (28)
Observe that the definition of $V_{N}(\mathbf{x})$ in (26) implies
$V_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi).$
differentiating on both sides with respect to $\xi$ gives
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}V_{N}(\mathbf{x}+\xi)=\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
Evaluation at $\xi=0$ and using (4.1) proves the assertion.
Figure 3: Finite-horizon and steady-state GNEP with GNE-KKT system
4.2 Sensitivity analysis of the storage function
The next result translates a crucial observation made in the context of turnpike analysis of optimal control problems to GNEPs. In some sense, the next result provides a storage function counterpart to Theorem 6.
Theorem 7 (Storage gradient).
Consider GNEP (2), its steady-state counterpart (10), and let Assumptions 23 hold. Then, the following statements hold:
  1. (i)
    Any primal-dual solution $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ of (25) constitutes a steady-state solution of (24a24c).
  2. (ii)
    Let $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v}=0,\bar{\nu}^{v})$ solve (25) such that $(\bar{x},\bar{u})\in\Pi\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ for some $\mathbf{x}\in\mathcal{X}_{0}$. Then, for any storage function $\Lambda$ differentiable at $\bar{x}$, it holds that
    $\sum_{v\in\mathcal{V}}\bar{\lambda}^{v}=-\nabla\Lambda(\bar{x}).$ (29)
Proof 4.2.
Part (i): Notice that at steady state we can neglect the initial condition for $x$ and the terminal condition for $\lambda^{v}$. Then, assertion (i) follows from direct comparison of (25) and (24).
Part (ii): The first sentence of part (ii) gives that the strict dissipation inequality (11) can be evaluated at $(\bar{x},\bar{u})$. Indeed, at $(\bar{x},\bar{u})$, the strict dissipation inequality (11) holds with equality. Differentiating it at $\bar{x}$ gives
$0=\ell_{x}-\left(f_{x}^{\top}-I\right)\nabla\Lambda(\bar{x})$ (30)
with $\ell$ from (11).
Next, we sum (25b) over $v\in\mathcal{V}$ and obtain
$0=\sum_{v\in\mathcal{V}}\ell_{x}^{v}+\left(f_{x}^{\top}-I\right)\lambda^{v}=\ell_{x}+\left(f_{x}^{\top}-I\right)\sum_{v\in\mathcal{V}}\bar{\lambda}^{v}$
where we use (11) and the fact that $\bar{\mu}^{v}=0$. Comparison with (30) directly gives the assertion.
Note that Theorem 7 (i) states that (25a25c) is a fixed point for (24a24c) but not for (2) as this would also require the boundary conditions (24d) to hold.
Moreover, the negative sign in (29) stems from the steady-state constraint $f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}=0$ in (10). If this constraint is equivalently written as $\bar{x}-f(\bar{x},\bar{u}^{v},\bar{u}^{-v})=0$, then corresponding Lagrange multiplier flips sign and the minus in (29) can be dropped.
If $\bar{\mu}^{v}\not=0$, there exist active per-agent constraints in the steady-state problem (10). The previous result can then be extended to
$\sum_{v\in\mathcal{V}}\bar{\lambda}^{v}+\left(f_{x}^{\top}-I\right)^{\dagger}g_{x}^{\top}\bar{\mu}^{v}=-\nabla\Lambda(\bar{x}),$
where $(\cdot)^{\dagger}$ denotes the Moore-Penrose inverse.
The above result shows two important links between the GNEP (2) and its steady-state counterpart (10): The optimality systems of both problems are closely related which resembles insights from optimal control theory [Trélat and Zuazua, 2015, Zanon and Faulwasser, 2018, Faulwasser and Grüne, 2022]. Moreover, there exists a strong link between the dual variables $\lambda^{v}$ and the storage function $\Lambda$. In the context of optimal control problems similar results have been shown in [Damm et al., 2014, Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018]. Yet, in the context of GNEPs this link is novel, as the dissipativity-based analysis of GNEPs is still in its infancy.
We conclude this section with a corollary that evaluates (29) at the turnpike state $x_{s}$.
Corollary 8.
Consider the setting of Theorem 7. If $\bar{x}=x_{s}$ we have that
$\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s}),$
where, for all $v\in\mathcal{V}$, $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ is a KKT point solving (25).
If the GENP exhibits a turnpike at $(x_{s},u_{s})$ and horizon $N$ is sufficiently large, we arrive at a crucial link between the value function (26) and any differentiable storage function
$\nabla V_{N}(x_{s})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0}\approx\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s})$
or, for all $v\in\mathcal{V}$, $\lambda^{v}_{0}\approx\lambda_{s}^{v}$.
4.3 Turnpike leaving arc and linear end penalties
A well-known aspect of the turnpike property is the characteristic leaving arc, where state and input trajectories diverge from the turnpike in the final part of the considered time horizon, cf. Figure 1. While the formal Definition 4 does not require that a turnpike solution has a leaving arc, the early works [Dorfman et al., 1958, McKenzie, 1976] and many later results [Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] focus on it.
In the context of optimal control, the presence of a leaving arc arises when the values of the states $x$ and the adjoints $\lambda$ close to the turnpike differs substantially from their values at the end of the horizon. In optimal control there are three mechanism which generate leaving arcs [Faulwasser and Grüne, 2022]:
  • Terminal constraints on the state variable such that $x_{N}=x^{N}\not=x_{s}$. Naturally this can be generalized to $x_{N}\in\mathbb{X}_{N}$.
  • Terminal stage costs $V_{\mathrm{f}}(x)$ which (without any active constraints at $k=N$) imply that the transversality condition $\lambda_{N}=\nabla V_{\mathrm{f}}(x)$.
  • Or the presence of state constraints which render it cheap to let a system drift towards the constraint boundary at the end of the horizon.
Interestingly, the same leaving arc phenomenon can be observed in GNEPs, see, e.g., the supply chain example [Hall et al., 2024a]. Moreover, the optimality interpretation of GNEPs—and in particular the structure of the KKT conditions, which is very similar to classic optimal control—suggest that the same mechanisms generate a leaving arc.
Inspired by the OCP analysis in [Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018], the next result shows how to suppress leaving arcs.
Proposition 9 (GNEPs without leaving arcs).
Consider the GNEP (2) and let Assumptions 23 hold. Let $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ be a KKT point solving (25). Consider the following two statements:
  1. (i)
    For all $v\in\mathcal{V}$, in (2) the per-agent cost function is of the form
    $\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})+V_{\mathrm{f}}^{v}(x_{N})$
    and
    $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}.$
  2. (ii)
    For all $v\in\mathcal{V}$, in (2) the point-wise terminal constraint $x_{N}=x_{s}$ is considered.
Then, for all $N\in\mathbb{N}_{>0}$, the constant trajectory $(x,u)\equiv(x_{s},u_{s})$ satisfies
$(x_{s},u_{s})\equiv:(x,u)=(x^{*},u^{*})\in\mathcal{S}_{N}^{GNE}(x_{s}).$
Proof 4.3.
As per Theorem 7, part (i,) the steady-state primal-dual tuple $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ from (25) is a steady state of (24a24c). Moreover, using $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ to construct an $N$-step constant trajectory also the complementary slackness and the dual feasibility conditions (24e-24f) are satisfied as they correspond point-wise in time to their counterparts in (25).
It remains to analyze (24d). For case (i), notice that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ in (23) has an additional term
$V_{\mathrm{f}}^{v}(x_{N})$
which, by standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda_{N}=\nabla V_{\mathrm{f}}^{v}(x_{N})+g_{x}^{\top}\mu_{N}^{v}.$ (24d-i)
Recall that the assertion sets $\mathbf{x}=x_{s}$. From the condition that $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$ it then immediately follows that the $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies $x_{0}=x_{s}$ and (24d-i).
For case (ii), observe that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ now includes the additional term
$\sigma^{\top}(x_{N}-x_{s}),$
where $\sigma\in\mathbb{R}^{n}$ is the additional multiplier for the point-wise terminal constraint. Once more, standard KKT analysis leads to the modified boundary condition
$x_{0}=\mathbf{x}\quad\text{and}\quad x_{N}=x_{s}.$ (24d-ii)
The $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},$ $\mu_{s}^{v},\nu_{s}^{v})$ satisfies these boundary conditions as $\mathbf{x}=x_{s}$.
The preceding analysis has shown that the constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v})$ satisfies the KKT conditions (24). Theorem 5 then directly gives the assertion. This finishes the proof.
A first connection between the KKT system and the role of the terminal constraint on the dual variable was drawn in [Hall et al., 2024b, Prop. 3] and the proof therein. Specifically, it was demonstrated that with a terminal constraint $x_{N}=x_{s}$ and additional assumptions on the cost, the steady-state GNEP $(x_{s},u_{s})$ is the unique fixed point of the finite-horizon GNEP-KKT system and the unique equilibrium point of a closed-loop system resulting from a receding-horizon implementation of the GNE.
Corollary 10 (Linear end penalty).
Let the conditions of Proposition 9 hold and suppose that $\mu_{s}^{v}=0$ for all $v\in\mathcal{V}$. Then
$V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v},\quad\forall v\in\mathcal{V}$
gives that, for any horizon $N\in\mathbb{N}_{>0}$, the constant trajectory $(x,u)\equiv(x_{s},u_{s})$ satisfies
$(x_{s},u_{s})\equiv:(x,u)=(x^{*},u^{*})\in\mathcal{S}_{N}^{GNE}(x_{s}).$
Indeed, quite similarly to the optimal control setting [Zanon and Faulwasser, 2018], also for GNEPs (2), the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ admits a primal and a dual interpretation:
  • It directly specifies the boundary or terminal constraint on the dual variable $\lambda^{v}_{N}$. Interestingly, the point-wise boundary condition on $\lambda^{v}_{N}$ is an optimality conditions but it does not imply any reachability requirement of the actual primal state dynamics $x^{+}=f(x,u)$.
  • Via a converse telescopic sum argument $V_{\mathrm{f}}^{v}(x)=\lambda_{s}^{\top}x$ can be pulled into the stage cost
    $\tilde{\ell}^{v}(x,u^{v},u^{-v})=\ell^{v}(x,u^{v},u^{-v})+\lambda_{s}^{\top}\left(f(x,u^{v},u^{-v})-x\right).$
    This trick is called rotating the stage cost in economic MPC [Diehl et al., 2011, Angeli et al., 2012]. The linear rotation with $\lambda_{s}$ implies that the steady constraint in (10) is only weakly active. This geometric modification of the stage cost with $\lambda_{s}$ has been called gradient correction in economic MPC, cf. [Zanon and Faulwasser, 2018].
4.4 Learning the linear end penalty
One may ask about the pros and cons of the different approaches that allow suppressing the leaving arc of turnpike solutions. Indeed, the point-wise terminal constraint $x_{N}=x_{s}$ requires solving the steady-state GNEP (10) as does the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$. In settings where the cost function changes occasionally—e.g. due to rare parametric updates (prices, etc.)—it is desirable to avoid the extra computational effort. Also, the primal terminal constraint $x_{N}=x_{s}$ jeopardize feasibility in case of short horizons.
In the context of suppressing turnpike leaving arcs in OCPs, [Faulwasser and Zanon, 2018, Sec. 4] proposed an adaptive strategy for the terminal penalty.
Algorithm 1 summarizes how this strategy may be transferred to GNEPs. In Step 2 the GNEP (2) is solved considering the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, where the initialization $p_{0}=0$ is used. Then, based on our turnpike insights, for all $v\in\mathcal{V}$, an approximation of the turnpike dual variable $\lambda^{v}_{s}$ is obtained by $p_{i}^{v}=\lambda^{v}(\frac{N}{2})$ in Step 5. The GNEP is resolved using the updated end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$. The algorithm ends if either the change in $\|p_{i}^{v}-p_{i+1}^{v}\|$ becomes small or an iteration limit is reached. Put differently, the algorithm builds upon the observation that whenever the states and inputs are close to their turnpike values also the dual variables stay close to their turnpike values. It is worth remarking that each agent can learn its specific linear end penalty. One may even learn this end penalty for some agent while others still exhibit leaving arcs.
While the formal convergence analysis is beyond the scope of the present paper, the next section will present numerical results demonstrating the efficacy of the proposed approach.
Algorithm 1 Learning $\lambda_{s}^{v}$ [Faulwasser and Zanon, 2018]
1:GNEP (2) with horizon $N=2\hat{N}$, initial condition $\mathbf{x}$, and $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, $i=0,p_{0}=0$, $i_{max}\in\mathbb{N}$, $\Delta_{0}=\infty$, $\varepsilon>0$
[…]
3:while $i\leq i_{max}$ OR $\Delta_{i}\leq\varepsilon$ do
4:  Solve GNEP (2) with $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$
5:  $i\leftarrow i+1$
6:  for all $v\in\mathcal{V}$ do
7:   $p_{i}^{v}\leftarrow\lambda^{v}(\frac{N}{2})$
8:  end for
9:  $\Delta_{i}\leftarrow\sum_{v\in\mathcal{V}}\|p^{v}_{i}-p^{v}_{i-1}\|$
10:end while
11:return $\hat{\lambda}_{s}^{v}=p_{i}^{v},v\in\mathcal{V}$
5 Simulation Study
We consider a simple example of the GNEP in (2) in which we have coupled LTI dynamics, coupled costs, and constraints $\forall v\in\mathcal{V}=\{1,2\}:$
$\left\{\begin{array}[]{r l}\displaystyle\min_{u^{v},x}&\displaystyle\sum_{k=0}^{N-1}\displaystyle u_{k}^{v}(\sum_{j\in\mathcal{V}}R^{v,j}u_{k}^{j})+\|x_{k}-x^{\text{ref}}\|_{Q^{v}}^{2}\\ \textrm{s.t.}&x_{k+1}=Ax_{k}+\displaystyle\sum_{j\in\mathcal{V}}B^{j}u^{j},\hskip 1.00006ptk\in\mathbb{Z}_{N}\\ &-2\leq u_{k}^{v}\leq 2,\hskip 46.00012ptk\in\mathbb{Z}_{N}\\ &-2\leq\displaystyle\sum_{j\in\mathcal{V}}u_{k}^{j}\leq 2,\hskip 28.99997ptk\in\mathbb{Z}_{N}\\ &-1\leq x_{k}\leq 1,x_{0}=1.\hskip 13.00005ptk\in\mathbb{Z}_{N+1}.\end{array}\right$. (31)
The parameter values are $A=1.5$, $B^{1}=1$, $B^{2}=2$, $R^{1,1}=R^{1,2}=4$, $R^{2,2}=R^{2,1}=5$ and state weights $Q^{1}=1,Q^{2}=2$. The reference state is chosen as $x^{\text{ref}}=0.3$. Assumptions 2 and 3 are fulfilled by construction.
The game pairs $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ for different horizon lengths are shown in Figure 4(a) and the corresponding trajectory of dual variables $\lambda^{v}$ in Figure 4(b). We clearly see the characteristic turnpike phenomenon of both input and state trajectories converging to the steady-state GNE $(x_{s},u_{s})$ and diverging in the final time steps.
Figure 4: Open-loop GNE trajectories of (31) without terminal penalty.
As outlined in Section 4.3 we apply a terminal penalty to the cost function of each agent which becomes $J^{v}(x,u^{v},u^{-v})+(\lambda_{s}^{v})^{\top}x_{N}$ the resulting trajectories are presented in Figure 5. Notice that with the terminal penalty, the state and input trajectories converge to the turnpike and remain there until the end of the horizon. In Figure 6, we demonstrate that in the very simple GNE problem (31) the penalty learning Algorithm 1 suppresses the leaving arc substantially even after just one iteration, i.e., $\lambda^{v}_{k}$ is close to the steady-state multipliers $\lambda^{v}_{s}$ at the midpoint of the trajectory.
Remark 4.
Since $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and the steady-state GNE in (10) are not necessarily unique, the terminal penalty method requires verifying that $\lambda_{s}$ corresponds to the global turnpike. Local turnpike phenomenon in OCPs have been studied in [Krügel et al., 2023], but this is left for future work in the GNE setting.
Figure 5: Open-loop GNE trajectories of (31) with a $\lambda_{s}\,x_{N}$ penalty.
Figure 6: Open-loop GNE trajectories of (31) with a learned terminal penalty following Algorithm 1 and stopping after one iteration.
6 Conclusion
In this manuscript we study GNE problems and derive system-theoretic insights into their open-loop trajectories based on dissipativity analysis. We extend the close relation between dissipativity and turnpike behaviors known in optimal control to dynamic generalized Nash games. Our simulation results show that the linear end penalties designed based on the optimality conditions successfully suppress the leaving arc of the turnpike even in the coupled, non-cooperative, game-theoretic setting.
In the future, we will investigate adaptive end penalties which would allow to learn $\lambda_{s}^{v}$ by extracting it from open-loop predictions while the game is being played rather than having to solve the steady-state GNE problem in advance. Further, we will study closed-loop stability based on turnpike results when GNEs are applied in a receding-horizon fashion as in MPC.

### v3

@@
-System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity
+System-Theoretic Analysis of Dynamic Generalized Nash Equilibria – Turnpikes and Dissipativity
 Abstract
-Generalized Nash equilibria are used in multi-agent control applications to model strategic interactions between agents that are coupled in the cost, dynamics, and constraints.
-We study the properties of open-loop GNE trajectories from a system-theoretic perspective.
+Generalized Nash equilibria are used in multi-agent control applications to model strategic interactions between agents that are coupled in the cost, dynamics, and constraints, and provide the foundations for game-theoretic MPC (Receding Horizon Games).
+We study properties of finite-horizon dynamic GNE trajectories from a system-theoretic perspective.
 We show how strict dissipativity generates the turnpike phenomenon in GNE solutions.
@@
 We derive conditions under which the steady-state GNE is the optimal operating point and, using a game value function, we give a local characterization of the geometry of storage functions.
-Finally, we design linear terminal penalties that ensure GNE open-loop trajectories converge to and remain at the steady-state GNE.
-These connections provide the foundation for future system-theoretic analysis of GNEs similar to those existing in optimal control.
+Finally, we design linear terminal penalties that ensure dynamic GNE trajectories applied in open-loop converge to and remain at the steady-state GNE.
+These connections provide the foundation for future system-theoretic analysis of GNEs similar to those existing in optimal control as well as for recursive feasibility and closed-loop stability results of game-theoretic MPC.
 keywords:
-dynamic games, optimal control, Nash equilibrium problems, multi-agent systems, dissipativity, turnpike property, non-cooperative game theory
+dynamic games, optimal control, generalized Nash equilibrium problems, multi-agent systems, dissipativity, turnpike property, non-cooperative game theory
 thanks: The authors are with the Automatic Control Lab, ETH Zürich, Physikstrasse 3, 8092 Zürich, Switzerland and with the Institute of Control Systems, Hamburg University of Technology, Harburger Schloßstraße 22a, 21079 Hamburg, Germany.
@@
 1 Introduction
-Control is at the core of many multi-agent applications managing interactions between strategic agents with conflicting objectives and coupled action spaces.
-Generalized Nash equilibria (GNEs) have emerged as a promising solution concept at the intersection of game theory and optimal control [Facchinei and Kanzow, 2009].
+Control is at the core of many multi-agent applications managing interactions between strategic agents with conflicting objectives, dynamics, and coupled action spaces.
+Generalized Nash equilibria (GNEs) have emerged as a promising solution concept at the intersection of game theory and optimal control [24].
 They naturally enable to define per-agent optimal control problems that are coupled, in a game-theoretic sense, in the objective, constraints, and dynamics.
-GNEs have been applied to resource allocation problems (e.g., energy [Atzeni et al., 2012, Hall et al., 2022], transportation [Bassanini et al., 2002], telecommunications [Pavel, 2012]) and competitive settings (e.g., autonomous driving [Dreves and Gerdts, 2017, Le Cleac’h et al., 2022], supply chains [Hall et al., 2024a]).
-The prospect for applications has catalyzed extensive research on efficient GNE-seeking algorithms and their convergence guarantees [Facchinei et al., 2010, Belgioioso et al., 2022].
-The resulting solution sets have been studied in detail [Kulkarni and Shanbhag, 2012, Nabetani et al., 2009], as well as their efficiency [Kulkarni, 2019] and fairness properties [Hall et al., 2025].
+GNEs have been applied to resource allocation problems in dynamic settings (e.g., energy [4, 36], transportation [5], telecommunications [51]) and competitive dynamic settings (e.g., autonomous driving [21, 43], supply chains [38]).
+Applications in real-time control of dynamic multi-agent systems catalyzed extensive research on efficient GNE-seeking algorithms and their convergence guarantees [32, 6].
+The resulting solution sets have been studied in detail [41, 48], as well as their efficiency [42] and fairness properties [37].
 A fundamental step towards safe deployment of controllers is the characterization of their closed-loop properties.
 Yet, the behavior of finite-horizon trajectories resulting from GNEs remains poorly understood, even in the open-loop case.
-Previous works have applied dissipativity techniques to analyze the feedback mapping [Hall et al., 2024b] and drawn connections to dynamic variational inequalities [Benenati and Grammatico, 2024].
-Despite this progress finite-horizon, GNE trajectories lack the rich system-theoretic characterizations developed for optimal control problems over the past decades, cf.
-[Willems, 1971, Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] and many others.
-In classic optimal control, necessary first-order conditions are given by the maximum principle [Boltyanskii et al., 1960] while dynamic programming provides necessary and sufficient conditions [Bellman, 1954].
-Moreover, the turnpike phenomenon is a similarity property of parametric optimal control problems with respect to varying initial conditions and horizon lengths.
-In essence, the phenomenon implies that for different initial conditions and horizon lengths the optimal trajectories approach the neighborhood of an optimal operating point (entry arc), remaining there (turnpike arc) and then diverge from it towards the end of the horizon (leaving arc).
-First observations date back to [Ramsey, 1928] and [von Neumann, 1938], while the name turnpike was coined by [Dorfman et al., 1958].
-We refer to [McKenzie, 1976] and [Faulwasser and Grüne, 2022] for classic and modern literature overviews.
-A recent key development is the analysis of the links of the turnpike phenomenon with dissipativity properties [Grüne and Müller, 2016, Faulwasser et al., 2017], which is in turn linked to the dissipativity route to infinite-horizon optimal control which originated with [Willems, 1971], see [Faulwasser and Kellett, 2021] for recent nonlinear extensions.
-For results on turnpike analysis exploiting the structures of the maximum principle we refer to [Trélat and Zuazua, 2015].
-In applications such as energy management, the turnpike effect or end-of-horizon effects lead, for instance, to myopic discharging of batteries [Hall et al., 2022, Vincent et al., 2020].
-Interestingly, turnpike phenomena have also been observed in games in the Economics literature as early as the 1980s [Fershtman and Muller, 1986] and was later treated from a theoretical perspective for infinite-horizon open-loop games in continuous time [Carlson and Haurie, 1995] and discrete time [Carlson and Haurie, 1996] focusing on existence, uniqueness of equilibria and convergence to the turnpike.
-Extensions to differential games with coupled state constraints followed [Carlson and Haurie, 2000].
-The turnpike property in games has recently received renewed attention.
-Turnpike properties for stochastic differential LQ games are studied in [Li et al., 2025] for the zero-sum setting and in [Cohen and Jian, 2025] for the non-zero sum setting.
-Multiple works derive turnpike properties for mean field games under a large population assumption [Cirant and Porretta, 2021, Carmona and Zeng, 2024, Ersland et al., 2025, Fedorov, 2025].
-Further, turnpikes were observed in competitive supply chain games [Hall et al., 2024a].
-This paper builds a bridge between the dissipativity route to turnpikes in optimal control and dynamic games.
-We establish the turnpike property for discrete-time finite-horizon GNE problems and study its implications on dissipativity of the GNE with respect to the steady-state GNE.
-Specifically, we make the following four contributions:
+Previous works have applied dissipativity techniques to analyze the feedback mapping [35] and drawn connections to dynamic variational inequalities [8].
+Despite this progress, finite-horizon GNE trajectories lack the rich system-theoretic characterizations developed for optimal control problems over the past decades, cf.
+[57, 2, 54] and many others.
+Figure 1: Schematic of a dynamic GNE state trajectory $x^{*}$ exhibiting the turnpike property at the steady-state GNE $x_{s}$ for different horizon lengths.
+In this context, the turnpike phenomenon arises as a property of solutions to an optimal control problem (OCP) solved for different initial conditions and varying horizon lengths.
+The phenomenon is characterized by the solutions clustering near one common steady state which is referred to as the turnpike, see Figure 1.
+It is a similarity property of parametric OCPs and thus is a crucial element in the closed-loop analysis of OCPs applied in a receding-horizon fashion, i.e., in model predictive control (MPC).
+First observations date back to [52] and [56], while the notion turnpike was coined by [20].
+We refer to [45] and [28] for classic and modern literature overviews.
+A recent key development is the uncovering of the close connection between the turnpike phenomenon and dissipativity properties [33, 27], which is in turn linked to the dissipativity route to infinite-horizon optimal control which originated with [57], see [26] for recent nonlinear extensions.
+The link between turnpike and dissipativity enabled major breakthroughs in the closed-loop stability analysis of nonlinear and economic MPC, see, e.g., [34].
+In applications such as energy management, the turnpike effect or end-of-horizon effect leads, for instance, to myopic discharging of batteries [36, 55].
+Interestingly, turnpike phenomena have also been observed in games in the Economics literature as early as the 1980s [31] and were later treated from a theoretical perspective for infinite-horizon open-loop games in continuous time [11] and discrete time [12] focusing on existence, uniqueness of equilibria and convergence to the turnpike.
+Extensions to differential games with coupled state constraints followed [13].
+Turnpike properties in games have recently received renewed attention: The specialized class of stochastic differential LQ games with continuous-time dynamics are studied in [44] for the zero-sum setting and in [17] for the non-zero sum setting.
+Multiple works derive turnpike properties for mean field games under a large population assumption [16, 14, 23, 30].
+Further, turnpikes were observed in competitive dynamic supply chains [38].
+This paper lays the foundations for an overarching theory of turnpike and dissipativity analysis in noncooperative games for a general class of cost functions, constraints, and discrete-time dynamics.
+We build a bridge between the dissipativity route to turnpikes in optimal control and dynamic games.
+Specifically, we establish the turnpike property for finite-horizon discrete-time dynamic GNE problems and study its implications on dissipativity of the GNE with respect to the steady-state GNE.
+Overall, we make the following four contributions:
   1. 1.
-    Demonstrating the crucial structural link between turnpike properties in parametric OCPs and parametric GNEs providing a foundation for system-theoretic analysis of GNEs.
+    We demonstrate the crucial structural link between turnpike properties in parametric OCPs and parametric GNEs providing a foundation for system-theoretic analysis of GNEs.
   2. 2.
-    Establishing that strict dissipativity implies the turnpike property, also in GNEs.
+    We establish that strict dissipativity implies the turnpike property, also in GNEs.
 Further, we derive a converse turnpike result, connecting the turnpike of the open-loop trajectory to dissipativity with respect to the steady-state GNE.
   3. 3.
-    Providing an optimality-like interpretation of GNEs by considering a game value function and studying its local gradient structure.
+    We provide an optimality-like interpretation of GNEs by considering a game value function and studying its local gradient structure.
 Additionally, we show that the storage function’s gradient equals the sum of agents’ dual multipliers at the steady-state GNE.
   4. 4.
-    Designing mechanisms to suppress the leaving arc in GNE open-loop trajectories, using per-agent linear terminal penalties to ensure trajectories converge to and remain at the optimal operating point.
+    We design mechanisms to suppress the leaving arc in GNE open-loop trajectories, using per-agent linear terminal penalties to ensure trajectories converge to and remain at the steady-state GNE.
 The remainder of this paper is organized as follows.
@@
 […]
-2 Problem Setting: Finite-Horizon Generalized Nash Equilibrium Problem
-A group of self-interested agents $v\in\mathcal{V}:=\{1,\dots,M\}$ controls the following shared dynamics
-$x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v})$ (1)
-for a compact set of initial conditions $\mathbf{x}\in\mathbb{X}_{0}$.
-Each agent minimizes their objective function over a prediction horizon $N$ with states and inputs coupled to other agents.
-This constitutes a finite-horizon Generalized Nash Equilibrium problem (GNEP) as follows
+2 Finite-Horizon Generalized Nash Equilibrium Problems
+Consider a nonlinear discrete-time system
+$x_{k+1}=f(x_{k},u_{k}),\quad x_{0}=\mathbf{x}$
+with state $x_{k}\in\mathbb{R}^{n_{x}}$, control input $u_{k}\in\mathbb{R}^{n_{u}}$, and initial condition $\mathbf{x}$ which lies in a compact set $\mathbf{x}\in\mathbb{X}_{0}$.
+Suppose the input is not computed by a central entity but is instead a stacked vector $u_{k}=[(u_{k}^{1})^{\top},\ldots,(u_{k}^{M})^{\top}]^{\top}$, where each component $u_{k}^{v}$ is decided by a self-interested agent $v\in\mathcal{V}:=\{1,\dots,M\}$.
+Each agent aims to steer the shared dynamics in its own favor
+$x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v}),\quad x_{0}=\mathbf{x}.$ (1)
+Specifically, each agent $v$ minimizes its accumulated stage cost $\ell^{v}$ over a prediction horizon $N$ subject to the shared dynamics (1).
+In addition to this dynamic coupling, agents influence each other’s outcomes through coupling in their cost and constraints, while still being subject to individual local constraints.
+Formalizing this gives rise to a finite-horizon dynamic Generalized Nash Equilibrium Problem (GNEP) of the form
 $\displaystyle\min_{u^{v},\,x}$ $\;\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})$ (2a)
@@
 $g(x_{k},u_{k}^{v},u^{-v}_{k})\leq 0,\hskip 25.00003ptk\in\mathbb{Z}_{N}$ (2c)
-$G^{v}\,u_{k}^{v}\leq h^{v},\hskip 54.00002ptk\in\mathbb{Z}_{N}$ (2d)
-$\;x_{0}=\mathbf{x}.$ (2e)
-where $\mathbf{x}$ is the initial condition of system (1). We define the following per-agent and global action sets
+$h^{v}(u_{k}^{v})\leq 0,\hskip 56.00014ptk\in\mathbb{Z}_{N}$ (2d)
+$\;x_{0}=\mathbf{x},$ (2e)
+where (2d) are nonlinear local constraints and (2c) nonlinear coupled constraints.
+For each choice of the initial condition $\mathbf{x}$ and the horizon $N$, the GNEP (2) generates finite-horizon GNE trajectories.
+Hence, we view the GNEP (2) as a parametric game-theoretic decision problem.
+Note that—for now—we make no assumptions on the functions in (2) but will add them when our theoretical results require them.
+We define the following per-agent and global action sets
 […]
 […]
-and the cumulative cost
-$J_{N}^{v}(x,u):=\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}).$
-Decisions that jointly solve (2) are generalized Nash equilibria (GNE) [Facchinei and Pang, 2009, §2], as defined next.
+as well as $\mathcal{Z}_{\infty}(\mathbf{x})$ for the infinite horizon problem and the per-agent cumulative cost
+$J_{N}^{v}(x,u^{v},u^{-v}):=\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u^{-v}_{k}).$
+With these abstractions we can write (2) as follows:
+$\forall v\in\mathcal{V}:\;\left\{\begin{aligned} \;\min_{u^{v},x}\;&J^{v}_{N}(x,u^{v},u^{-v})\\ \operatorname{\ {s.}{t.}\ }&(x,u^{v})\in\mathcal{Z}_{N}^{v}(\mathbf{x},u^{-v}).\end{aligned}\right$.
+(5)
+Decisions that jointly solve (2) are called generalized Nash equilibria (GNE) [25, §2]. Intuitively, at a GNE no agent $v\in\mathcal{V}$ can reduce its cost by unilaterally changing its own decision.
 Definition 1 (Generalized Nash equilibrium).
-A joint decision $(x^{*},u^{*})$ is a GNE of (2) if $\forall v\in\mathcal{V}$
-$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})$
-$\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$. The corresponding solution set for fixed $N\in\mathbb{N}$ and $\mathbf{x}\in\mathbb{X}_{0}$ is denoted as
+A joint decision $(x^{*},u^{*})\in\mathcal{Z}_{N}(\mathbf{x})$ is a GNE of (2) if
+$\forall v\in\mathcal{V}:\,J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})$
+$\forall(x,u^{v})\in\mathcal{Z}_{N}(\mathbf{x},u^{-v*})$. The corresponding solution set for fixed $N\in\mathbb{N}$ and $\mathbf{x}\in\mathbb{X}_{0}$ is denoted as
 $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\subset\mathbb{R}^{(N+1)n_{x}+Nn_{u}}.$ (6)
-Intuitively, a GNE is a joint decision for which no agent $v\in\mathcal{V}$ can reduce its cost by unilaterally changing its own decision.
-For the sake of concise exposition and in analogy to optimal control concepts, we refer to $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ as a GNE pair or an optimal game pair.
-Moreover, with slight abuse of notation we write
-$(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$
-to express that there exist $k\in\mathbb{Z}_{N}$ and $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ such that
-$(x_{k},u_{k})=(x_{k}^{*},u_{k}^{*})\quad\text{with}\quad(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}).$
-We want to study the convergence of the finite-horizon GNE trajectory with respect to a steady-state point. Thus, we introduce the game-theoretic equilibrium of the steady-state game next.
+In analogy to optimal control concepts, we refer to $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ as a GNE pair or an optimal game pair.
+As a shorthand, the point-wise in time projection of $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ onto the states is written as $\mathbb{X}_{N}(\mathbb{X}_{0})\subseteq\mathbb{R}^{n_{x}}$.
+That is, if $\tilde{x}\in\mathbb{X}_{N}(\mathbb{X}_{0})$ then there exist at least one game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}),\mathbf{x}\in\mathbb{X}_{0}$ such that the state trajectory $x^{*}$ passes through $\tilde{x}$ at least once.
+Similarly $X_{\infty}(\mathbb{X}_{0})$ refers to the infinite horizon problem.
+Henceforth, we study convergence or clustering of finite-horizon GNE trajectories close to a two-fold equilibrium: a point that is both (i) a steady-state of (1), i.e., $\bar{x}=f(\bar{x},\bar{u}^{v},\bar{u}^{-v})$; and (ii) a strategic (decision) equilibrium of the one-step GNEP in (2) such that no agent benefits from deviating in the next time step.
+This is formalized in the following steady-state GNEP.
 Definition 2 (Steady-state GNE).
-The pair $(x_{s},u_{s})$ is called a steady-state GNE if it solves the steady-state GNEP
+The pair $(x_{s},u_{s})$ is called a steady-state GNE if it solves
 […]
-with the corresponding solution set $\mathcal{S}^{\text{\tiny GNE}}_{s}\subset\mathbb{R}^{n_{x}+n_{u}}.$
-We analyze properties of the game pair of the GNE in (2) provided a solution exists.
+with the corresponding solution set $\mathcal{S}^{\text{\tiny GNE}}_{s}\subset\mathbb{R}^{n_{x}+n_{u}}.$ Problem (11) is called a steady-state GNEP.
+We analyze properties of the GNE in (2) and (11) provided a solution exists.
 Thus, in the following we assume that $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $\mathcal{S}^{\text{\tiny GNE}}_{s}$ are nonempty $\forall\mathbf{x}\in\mathbb{X}_{0}$.
-Remark 1 (Existence & computation).
+Remark 1 (Existence and computation).
 The theoretical results which follow hold for the nonlinear GNE problem in (2).
-Existence of a GNE has been established under a variety of assumptions as summarized in [Dutang, 2013].
-Efficient algorithms to compute solutions of nonlinear GNEPs are a topic of ongoing research, one example is given in [Le Cleac’h et al., 2022].
-3 Dissipativity and Turnpike Relations for GNEPs
+Existence of a GNE has been established under a variety of assumptions, as summarized in [22].
+One approach is to assume that $\forall v\in\mathcal{V},\;\forall\mathbf{x}\in\mathbb{X}_{0},\;\forall u^{-v}\in\mathcal{U}^{-v}:=\Pi_{j\in\mathcal{V}\backslash v}\mathcal{U}^{j}\$the set $\mathcal{Z}_{N}^{v}(\mathbf{x},u^{-v})$ and $\mathcal{U}^{v}$ is nonempty, compact and convex and the function $J^{v}_{N}(\cdot,\cdot,u^{-v})$ is convex in both arguments [50, Prop.
+12.7].
+Efficient algorithms to compute solutions of nonlinear GNEPs are a topic of ongoing research, one example is given in [43].
+3 Dissipativity and Turnpikes for GNEPs
 Turnpike and dissipativity properties are closely linked, providing system-theoretic insights for optimal control.
 In the following, we demonstrate that this link also exists for the GNE problem in (2).
-We first introduce a performance measure for the population of agents as follows
+Consider the social welfare as a joint performance measure for the population of agents
 $\ell(x_{k},u_{k}):=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})$ (12)
-and
-$J_{N}(x,u):=\sum_{k=0}^{N-1}\ell(x_{k},u_{k})=\sum_{k=0}^{N-1}\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$
-While using the sum of costs may seem unusual from a game-theoretic perspective, [Benenati and Grammatico, 2024] show that it can serve as a Lyapunov function for LQ games with shared constraints under suitable controllability assumptions.
-This motivates our choice in (11) and simplifies the subsequent theoretical analysis.
-3.1 Dissipativity of GNEPs
-Based on the joint performance measure, we can make a strict dissipativity assumption analogous to the ones existing in optimal control.
+and introduce the shorthand
+$J_{N}(x,u):=\sum_{k=0}^{N-1}\ell(x_{k},u_{k}).$ (13)
+Remark 2 (GNEPs vs. OCPs).
+At this point it is fair to ask for the precise distinction between the GNE setting considered here and an optimal control point of view.
+Solving (5) yields a GNE, i.e., a joint decision from which no agent $v\in\mathcal{V}$ can reduce its cost unilaterally.
+In contrast, the optimal control counterpart of (5) reads
+$\;V^{\diamond}_{N}(\mathbf{x}):=\min_{u,x}\;J_{N}(x,u)\operatorname{\ {s.}{t.}\ }(x,u)\in\mathcal{Z}_{N}(\mathbf{x},u),$
+where all control actions are chosen with the sole objective of minimizing the cost for the entire agent population (13).
+Whenever necessary, we use superscript $\cdot^{*}$ for GNEP quantities and $\cdot^{\diamond}$ for OCP quantities.
+Take a simple unconstrained quadratic example with two agents and cost $J^{1}(u^{1},u^{2})=r^{1,1}(u^{1})^{2}+r^{1,2}u^{1}u^{2}+r^{1}_{\text{lin}}u^{1}$, and analogously for agent 2.
+Any equilibrium solution must satisfy
+$\left[\begin{smallmatrix}\frac{\partial J^{1}(u^{1},u^{2})}{\partial u^{1}}\\ \frac{\partial J^{2}(u^{1},u^{2})}{\partial u^{2}}\end{smallmatrix}\right]=\underbrace{\left[\begin{smallmatrix}2r^{1,1}&r^{1,2}\\ r^{2,1}&2r^{2,2}\end{smallmatrix}\right]}_{R^{*}}u+\left[\begin{smallmatrix}r^{1}_{\text{lin}}\\ r^{2}_{\text{lin}}\end{smallmatrix}\right]\stackrel{{\scriptstyle!}}{{=}}0$
+whereas $R^{\diamond}=\left[\begin{smallmatrix}2r^{1,1}&0.5(r^{1,2}+r^{2,1})\\ 0.5(r^{1,2}+r^{2,1})&2r^{2,2}\end{smallmatrix}\right]$ and thus $u^{*}$ and $u^{\diamond}$ diverge with diverging $r^{1,2}$ and $r^{2,1}$.
+It is well known that in noncooperative games where agents act selfishly, the equilibrium solution does not minimize the sum of costs, and thus differs from the OCP solution.
+This maximal efficiency loss is referred to as the price of anarchy (PoA); see [42] for an analysis in the context of GNEs.
+Formally,
+$\text{PoA}(\mathbf{x}):=\dfrac{\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})}{V^{\diamond}_{N}(\mathbf{x})}\geq 1.$ (14)
+Next, using the joint performance measure $J_{N}$ in (13), we introduce a strict dissipativity notion which follows concepts used in optimal control.
 Definition 3 (Strict dissipativity of GNEPs).
-Given a steady-state GNE $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$, the GNEP (2) is called strictly dissipative if for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ there exists a storage function $\Lambda:\mathbb{X}_{0}\to\mathbb{R}$ bounded from below and $\alpha_{\ell}\in\mathcal{K}$ such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$
+Given a steady-state GNE $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$, the GNEP (2) is called strictly dissipative with supply rate
+$s(x_{k},u_{k}):=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$
+if there exists a storage function $\Lambda:\mathbb{X}_{N}(\mathbb{X}_{0})\to\mathbb{R}$ bounded from below such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$
 $\Lambda(f(x_{k},u_{k}))-\Lambda(x_{k})\leq\\ -\alpha_{\ell}\left(\left\|\begin{matrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{matrix}\right\|\right)+s(x_{k},u_{k})$ (14)
-holds for each $(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
-Figure 1: Schematic of a state trajectory exhibiting the turnpike property for different horizon lengths.
-[Willems, 1972] introduced equivalent dissipativity characterizations via the required supply and the available storage.
-In discrete time the available storage characterization can be traced back to [Byrnes and Lin, 1994].
-For our further investigations, the latter one is very helpful.
-In terms of GNEPs, this well-known characterization reads as follows.
-Theorem 1 (Available storage of GNEP ).
-The GNEP (2) is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ and $\alpha_{\ell}\in\mathcal{K}_{\infty}$ if and only if
-$\Lambda_{a}(\mathbf{x})=$ $\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x^{*},u^{*})\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}\sum_{k=0}^{N-1}-s(x_{k}^{*},u_{k}^{*})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)<\infty.$
-Note that Definition 3 and Theorem 1 consider strict dissipativity only along game pairs resulting from the GNEP in (2).
-In optimal control, strict dissipativity is often assumed to hold along all feasible trajectories, see for instance [Grüne and Müller, 2016, Defn.
-2.1], while [Faulwasser et al., 2017] also restricted the crucial inequality to hold only along optimal trajectories.
-However, in a game-theoretic setting we need to restrict the definition as the following shows.
-Remark 2 (Why restrict dissipativity to $\mathcal{S}^{\text{\tiny GNE}}_{N}?$).
-Denote the optimal steady state associated with (11) as
+holds for some $\alpha_{\ell}\in\mathcal{K}$ and each point $(x_{k},u_{k})$ along game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
+If (14) holds for $\alpha_{\ell}\equiv 0$, the GNEP (2) is called dissipative.
+Since $\alpha_{\ell}$ is of class $\mathcal{K}$, inequality (14) is strict whenever $(x_{k},u_{k})\neq(x_{s},u_{s})$, hence the term strict dissipativity is used in the literature.
+A natural question is how to verify strict dissipativity of GNEPs.
+To this end, consider
+$\Lambda_{\alpha_{\ell}}(\mathbf{x}):=\displaystyle\hskip-5.69054pt\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x^{*},u^{*})\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}\hskip-2.84526pt\sum_{k=0}^{N-1}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*}).$ (15)
+Observe that, given the initial condition $\mathbf{x}$, this is a free end-time optimization problem over all game pairs $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
+The objective is to compute the maximum storage that can be extracted from the system, hence $\Lambda_{\alpha_{\ell}}$ is referred to as the available storage [58, 10].
+The next result extends the classical characterization of dissipativity via available storage [58] to dynamic GNEPs.
+Theorem 1 (Available storage of GNEPs & (14)).
+The GNEP (2) is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ and $\alpha_{\ell}\in\mathcal{K}$ (Definition 3) if and only if, for all $\mathbf{x}\in\mathbb{X}_{\infty}(\mathbb{X}_{0})$, the bound $\Lambda_{\alpha_{\ell}}(\mathbf{x})<\infty$ holds.
+Proof 3.1.
+The proof is given in Appendix A.
+Willems’ motivation for the introduction of dissipativity notions in systems and control has been the “generalization of Lyapunov functions to open systems, to systems with inputs and outputs” [59].
+From this perspective, the storage function is the amount of energy in an open system and the available storage the amount that can be extracted via inputs and outputs.
+Thus, Theorem 1 states that strict dissipativity, which loosely implies that solutions are attracted to $(x^{s},u^{s})$, can only hold if for all initial conditions the energy extractable along GNE trajectories is bounded; if it were unbounded, no such attraction could occur.
+In OCPs, strict dissipativity is often assumed to hold along all feasible trajectories [33, Defn.
+2.1], though [27] restrict it to optimal ones.
+As the next result shows, in the game-theoretic setting the restriction to GNE trajectories of (2) is indeed necessary.
+Lemma 2.
+(Restriction of domain of (14)) The strict dissipativity inequality (14) does not necessarily hold along all feasible trajectories $(x,u)\in\mathcal{Z}_{N}(\mathbf{x})$.
+Proof 3.2.
+We introduce the optimal control counterpart of the steady-state game in (11) which gives the steady-state minimizer
 $(x^{\diamond},u^{\diamond})=$ $\operatorname*{argmin}_{\bar{x},\bar{u}}\;\ell(\bar{x},\bar{u})\;$ (16)
-$\text{s.t.}\;\bar{x}=f(\bar{x},\bar{u}),\;g(\bar{x},\bar{u})\leq 0,\;G\,\bar{u}\leq h.$
-Substituting it into (11) yields
-$\ell(x_{s},u_{s})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{\diamond}-x_{s}\\ u^{\diamond}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x^{\diamond},u^{\diamond}),$
-where $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$.
-This inequality, however, holds only if $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$ which implies that the steady-state GNE (10) has to coincide with the minimizer of (12), and thus be a solution to an optimization problem.
-In turn, this would require the game to be potential [Monderer and Shapley, 1996] and the potential function to be the sum of agents’ costs [Slade, 1994, Prp.
+$\text{s.t.}\;\bar{x}=f(\bar{x},\bar{u}),\;g(\bar{x},\bar{u})\leq 0,\;h(\bar{u})\leq 0.$
+in which $\ell(\bar{x},\bar{u})=\displaystyle\sum_{v\in\mathcal{V}}\ell^{v}(\bar{x},\bar{u}^{v},u^{-v})$ as defined in (12) and $h(\cdot)$ stacks all local constraints.
+Consider the dissipativity inequality in (14)
+$\Lambda(f(x_{k},u_{k}))-\Lambda(x_{k})\leq s(x_{k},u_{k})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$
+which evaluated at $(x^{\diamond},u^{\diamond})$ for $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ as given in Definition 3 yields
+$\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{\diamond}-x_{s}\\ u^{\diamond}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x^{\diamond},u^{\diamond})-\ell(x_{s},u_{s})$ (17)
+where $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$ is the steady-state GNE.
+By construction, $(x^{\diamond},u^{\diamond})$ minimizes $\ell(\bar{x},\bar{u})$.
+Hence (17) holds only if $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$, i.e., the steady-state GNE must coincide with the minimizer of (16).
+This, however, does not hold in general for the GNEP in (2) which asserts our claim.
+Example 1 (When do GNEP and OCP coincide?).
+We now provide intuition for when the price of anarchy (14) equals one, focusing on the steady-state problem (11) for simplicity.
+Indeed, one can identify conditions when $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$ by comparing the first-order optimality conditions of (11) and (16); they coincide if
+$\frac{\partial\ell^{v}(x_{k},u_{k})}{\partial x_{k}}=\frac{\partial\ell(x_{k},u_{k})}{\partial x_{k}}$ (18a)
+$\frac{\partial\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}=\frac{\partial\ell(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}$ (18b)
+holds $\forall v\in\mathcal{V}$ and as $\ell(x_{k},u_{k})=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k})$ this would require for all agents’ coupled cost function parts to be the same up to a constant.
+This would correspond to a potential game [46] with the sum of costs as the potential function [53, Prp.
 1].
-More specifically, we see from the first-order optimality conditions that it would require
-$\frac{\ell^{v}(x_{k},u_{k})}{\partial x_{k}}=\frac{\ell(x_{k},u_{k})}{\partial x_{k}}$ (13a)
-$\frac{\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}=\frac{\ell(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}$ (13b)
-and as $\ell(x_{k},u_{k})=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k})$ this would impose that all agents’ cost functions are the same up to a constant which does not hold for the GNEPs we consider in (2).
-It is a well known phenomenon that selfish agent behavior in games leads to solutions that are suboptimal compared to centrally minimizing the sum of costs (Price of Anarchy).
-[Kulkarni, 2019] analyze these efficiency losses for GNEs.
-Thus, in Definition 3 we require the dissipation inequality to hold only for trajectories resulting from the GNEP in (2).
-The next result shows that strict dissipativity also provides an avenue to analyze which $(x_{S},u_{s})\in\mathcal{S}^{GNE}_{S}$ can appear in (11).
-Lemma 2.
-Let the strict dissipation inequality (11) hold for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$. Then
-$\sum_{v\in\mathcal{V}}\ell^{v}(x_{s},u_{s}^{v})=:\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u}):=\sum_{v\in\mathcal{V}}\ell^{(}\bar{x},\bar{u}^{v})$
-holds for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$.
-Proof 3.1.
-For contradiction, suppose there exists a better steady state solution $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$ with $\ell(x_{s},u_{s})>\ell(\bar{x},\bar{u})$.
-At steady state, the strict inequality (11) gives
-$0<\alpha_{\ell}\left(\left\|\begin{matrix}\bar{x}-x_{s}\\ \bar{u}-u_{s}\end{matrix}\right\|\right)\leq\ell(\bar{x},\bar{u})-\ell(x_{s},u_{s}).$
-Since $\alpha_{\ell}\in\mathcal{K}$, this allows writing $\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u})$ and thus shows the contradiction.
-The last result requires (11) to hold on (10).
-This is needed, as without further analysis it is not clear that for $x_{0}=x_{s}$ the GNE control action at time $k=0$ is such that $u^{*}_{0}=u_{s}$.
-Indeed, the control action $u_{0}^{*}$ depends on the specific design of (2) and the specific horizon $N$.
-Corollary 10 in Section 4 presents a result that ensures $u^{*}_{0}=u_{s}$ if $x_{0}=x_{s}$.
-We now formally introduce the considered turnpike property for our GNE problem in (2) with respect to the steady-state GNE.
+We do not impose this restriction in (2).
+We now formally introduce the considered turnpike property for our GNE problem in (2) with respect to the steady-state GNE (11).
 Definition 4 (Measure turnpike in GNEPs).
-The GNEP (2) exhibits the (measure) turnpike property at $(x_{s},u_{s})$ if for each $\varepsilon>0$ there exists $C>0$ such that $\forall N\in\mathbb{N},\mathbf{x}\in\mathbb{X}_{0}$ and for all game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ it holds that
-$Q_{\varepsilon}:=\#\{k\in\mathbb{Z}_{N}|\,\|x_{k}-x_{s}\|\leq\varepsilon\}\geq N-\frac{C}{\alpha_{\ell}(\varepsilon)}.$ (14)
+The GNEP (2) exhibits the (measure) turnpike property at $(x_{s},u_{s})$ if for each $\varepsilon>0$ there exists $C>0$ such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$, and for all game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ it holds that
+$Q_{\varepsilon}:=\#\left\{k\in\mathbb{Z}_{N}|\,\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\leq\varepsilon\right\}\geq N-\frac{C}{\alpha(\varepsilon)}$ (19)
+for some $\alpha\in\mathcal{K}$ and where #\# refers to the cardinality.
+Definition 4 provides a lower bound on the number of time steps the GNE trajectory spends within $\mathcal{B}_{\varepsilon}(x_{s})$, an $\varepsilon$-ball around the steady-state GNE.
+This bound grows with the horizon $N$, thus intuitively as $N$ increases, the trajectory spends more time in a neighborhood around $x_{s}$, while the time spent outside $\mathcal{B}_{\varepsilon}(x_{s})$ remains bounded independently of $N$.
+Assumption 1 (Cheap reachability).
+For all $\mathbf{x}\in\mathbb{X}_{0}$, there exists an infinite-horizon feasible pair $(x,u)\in\mathcal{Z}_{\infty}(\mathbf{x})$ such that, for some $\delta\in\mathbb{R}$ and all $N\in\mathbb{N}$ it holds that
+$J_{N}(x,u)\leq N\ell(x_{s},u_{s})+\delta.$
+It is worth noting that this assumption is readily verified if from all $\mathbf{x}\in\mathbb{X}_{0}$, the equilibrium point $x_{s}$ can be reached in a finite number of steps or if the Jacobian linearization of (1) at $(x_{s},u_{s})$ is stabilizable.
+Moreover, we note that the feasible $(x,u)$ in Assumption 1 does not need to be a game pair.
+Assumption 2 (Bounded price of anarchy).
+For all $\mathbf{x}\in\mathbb{X}_{0}$, the price of anarchy from (14) satisfies
+$\text{PoA}(\mathbf{x})=\frac{\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})}{V^{\diamond}_{N}(\mathbf{x})}\leq P<\infty$
+for $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $0<\nu\leq V^{\diamond}_{N}(\mathbf{x})\leq V<\infty$ holds.
 We next show that strict dissipativity of the GNEP with respect to $(x_{s},u_{s})$ implies the turnpike property for the input and state trajectories resulting from (2).
-The result is inspired by [Grüne, 2013, Thm.
+The result is inspired by [34, Thm.
 5.3] and its proof.
-Theorem 3 (Str. dissipativity $\Rightarrow$ turnpike).
-If, for every $x\in\mathbb{X}_{0}$, the GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ in the sense of Definition 3, then all GNEP solutions with $J_{N}(x,u)\leq N\ell(x_{s},u_{s})+\delta$ exhibit the turnpike property at $(x_{s},u_{s})$.
-Figure 2: Overview of implications between strict dissipativity, turnpike and optimal operation at the steady-state GNE.
-Proof 3.2.
-We define $C:=2\sup_{\mathbf{x}\in\mathbb{X}_{0}}|\Lambda(\mathbf{x})|<\infty$ and further we have that
-$\tilde{J}_{N}(x,u)$ $=J_{N}(x,u)+\Lambda(\mathbf{x})-\Lambda(x_{N})$ (15a)
-$\leq J_{N}(x,u)+C,$ (15b)
-$\leq N\,\ell(x_{s},u_{s})+(\delta+C)$ (15c)
+Theorem 3 (Strict dissipativity $\Rightarrow$ turnpike).
+Consider the GNEP (2) and let Assumptions 1 and 2 hold.
+Suppose that the GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ in the sense of Definition 3 and the storage is bounded on $\mathbb{X}_{N}(\mathbb{X}_{0})$.
+Then all GNEP solutions exhibit the (measure) turnpike property at $(x_{s},u_{s})$.
+Proof 3.3.
+The proof proceeds in two steps.
+First we show that Assumptions 1 and 2 allows to bound the performance of the population of agents along game pairs.
+Second we exploit the strict dissipation inequality to construct the asserted bound (19).
+Step 1:  Lemma 39 from Appendix B shows that Assumption 2 implies the linear performance bound $J_{N}(x^{*},u^{*})\leq V^{\diamond}_{N}(\mathbf{x})+\tilde{P}$ for some finite $\tilde{P}$.
+Since Assumption 1 relies on non-optimal trajectories, we have the bound
+$J_{N}(x^{*},u^{*})\leq N\ell(x_{s},u_{s})+\delta+\tilde{P}$
+which holds for all $N\in\mathbb{N},\mathbf{x}\in\mathbb{X}_{0}$ and $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
+Step 2:  Let $C:=2\sup_{\mathbf{x}\in\mathbb{X}_{0}}|\Lambda(\mathbf{x})|<\infty$. Then we have
+$\tilde{J}_{N}(x^{*},u^{*})$ $:=J_{N}(x^{*},u^{*})+\Lambda(\mathbf{x})-\Lambda(x_{N})$ (20a)
+$\leq J_{N}(x^{*},u^{*})+C,$ (20b)
+$\leq N\,\ell(x_{s},u_{s})+(\delta+\tilde{P}+C)$ (20c)
 Also due to strict dissipativity we have that $\forall(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}),\,\forall\mathbf{x}\in\mathbb{X}_{0}$:
-$\ell(x_{k},u_{k})+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))\geq$
-$\ell(x_{s},u_{s})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\$(16)
-For contradiction suppose that the trajectory does not fulfill the turnpike property and thus, $Q_{\varepsilon}<N-\frac{\delta+C}{\alpha_{\ell}{(\varepsilon})}$.
-This implies that there exists a number of $N-Q_{\varepsilon}>\frac{\delta+C}{\alpha_{\ell}(\varepsilon)}$ times steps in the set $\mathcal{N}\subseteq\{0,\dots,N-1\}$ for which the trajectory is more than $\varepsilon$ distance away from the steady state $x_{s}$, i.e., $\|x_{k}-x_{s}\|>\varepsilon,\;\forall k\in\mathcal{N}$.
-$\tilde{J}_{N}(x,u)$ $=\sum_{k=0}^{N-1}[\ell(x_{k},u_{k}))+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))]$
+$\ell(x^{*}_{k},u^{*}_{k})+\Lambda(x^{*}_{k})-\Lambda(f(x^{*}_{k},u^{*}_{k}))\geq$
+$\ell(x_{s},u_{s})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)\$(21)
+Next we construct the bound from (19) with $\alpha=\alpha_{\ell}$.
+For contradiction suppose that the trajectory does not fulfill the turnpike property and thus, $Q_{\varepsilon}<N-\frac{\delta+\tilde{C}}{\alpha_{\ell}{(\varepsilon})}$ with $\tilde{C}=C+\tilde{P}$.
+This implies that there exists a number of $N-Q_{\varepsilon}>\frac{\delta+\tilde{C}}{\alpha_{\ell}(\varepsilon)}$ times steps in the set $\mathcal{N}\subseteq\{0,\dots,N-1\}$ for which the trajectory is more than $\varepsilon$ distance away from the steady state $(x_{s},u_{s})$, i.e., $\|(x^{*}_{k},u^{*}_{k})-(x_{s},u_{s})\|>\varepsilon,\;\forall k\in\{0,\dots,N-1\}$.
+We obtain
+$\tilde{J}_{N}(x^{*},u^{*})$ $=\sum_{k=0}^{N-1}[\ell(x^{*}_{k},u^{*}_{k}))+\Lambda(x^{*}_{k})-\Lambda(f(x^{*}_{k},u^{*}_{k}))]$
 […]
-$=N\ell(x_{s},u_{s})+Q_{\varepsilon}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)]$
-$+\underbrace{(N-Q_{\varepsilon})\,\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)}_{\geq(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)}$
+$=N\ell(x_{s},u_{s})+Q_{\varepsilon}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)]$
+$+\underbrace{(N-Q_{\varepsilon})\,\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)}_{\geq(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)}$
 $\geq N\ell(x_{s},u_{s})+(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)$
-$>N\ell(x_{s},u_{s})+(\delta+C)$
+[…]
 which clearly contradicts (20), thus the assertion.
-Theorem 3 demonstrates that strict dissipativity of the GNEP (2) is not merely a theoretical property but gives direct insights into the behavior of the trajectory with respect to the steady-state GNE.
+The proof relies on Assumptions 1 and 2 together with strict dissipativity with bounded storage.
+Assumption 2, specific to the GNEP setting, requires that the overall GNEP performance is not arbitrarily worse than its OCP counterpart, which is arguably a mild condition.
+Assumption 1 is not specific to GNEPs as it is a system property.
+It can be verified via reachability arguments as done in the OCP turnpike literature [34].
+Similarly, boundedness of the storage function $\Lambda$ is standard in dissipativity-based turnpike analysis.
+Actually Theorem 1 shows that the available storage is always bounded from above and below, so assuming strict dissipativity with bounded storage is no stronger than assuming (14) alone.
+Finally, one might argue that the dissipativity assumption in itself is artificial.
+Yet, Theorem 3 provides an explanation for the turnpike phenomenon in GNEPs and it elegantly links the dynamic game trajectories to the steady-state GNE.
+Moreover, we will see in Theorem 4 that also converse statements can be derived: turnpike $\Rightarrow$ dissipativity.
 Remark 3 ( Turnpike with strict $x$ dissipativity).
-Observe that—similar to its optimal control counterpart—the proof of Theorem 3 does not rely on $(x,u)$-dissipativity but only requires $x$-dissipativity with $\alpha_{\ell}(\|x-x_{s}\|)$.
-However, having strictness in $x$ and $u$ helps to derive converse turnpike results.
+Observe that—similar to its optimal control counterpart—the proof of Theorem 3 does not rely on $(x,u)$-dissipativity with $\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ but only requires $x$-dissipativity with $\alpha_{\ell}(\|x_{k}-x_{s}\|)$.
+That is, the result of Theorem 3 can be slightly generalized to show a measure turnpike at $x_{s}$ only.
+However, having strictness in $x$ and $u$ will later help to derive converse turnpike results.
 3.1 A converse turnpike result for GNEPs
-The question arises if the converse also holds, i.e., whether the turnpike behavior in the open-loop trajectories implies strict dissipativity with respect to $(x_{s},u_{s})$.
 We recall the following “off-equilibrium” concept.
-Definition 5 ($\varepsilon$-GNE [Chen et al., 2021]).
-For a given $\varepsilon\geq 0$, $u^{*}$ is said to be an $\varepsilon$-generalized Nash equilibrium ($\varepsilon$-GNE) of (2) if $\forall v\in\mathcal{V}$ and $\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$ it holds:
-$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})+\varepsilon,\;$
-The solution set is denoted as $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})$.
-Particularly, when $\varepsilon=0$, $u^{*}$ is a GNE and $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})=\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
-We make the following assumption which ensures that $(x_{s},u_{s})$ is the unique GNE in the set of $\rho$-GNEs centred at $x_{s}$ , $\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})$, which minimizes the performance measure $J(x,u)$ defined for the population of agents.
-Assumption 1 (Local cost minimizer $(x_{s},u_{s})$).
-There exists a constant $\rho$ and $\alpha_{\rho}\in\mathcal{K}$ such that $\forall(x,u)\in\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})\cap\mathcal{Z}(x_{s})$ and $\forall k\in\mathbb{N}$ it holds:
-$\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x_{k},u_{k})-\ell(x_{s},u_{s}).$ (17)
+Definition 5 ($\rho$-GNE [15]).
+For a given $\rho\geq 0$, the pair $(u^{*},x^{*})$ is said to be an $\rho$-generalized Nash equilibrium ($\rho$-GNE) of (2) if $\forall v\in\mathcal{V}$ and $\forall(x,u^{v})\in\mathcal{Z}_{N}(\mathbf{x},u^{-v*})$ it holds:
+$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})+\rho,\;$
+The solution set is denoted as $\mathcal{S}^{\rho\text{-\tiny GNE}}_{N}(\mathbf{x})$.
+Particularly, when $\rho=0$, $(u^{*},x^{*})$ is a GNE and $\mathcal{S}^{\rho\text{-\tiny GNE}}_{N}(\mathbf{x})=\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
+The following assumption ensures that, among all $\rho$-GNEs centred at $x_{s}$, the pair $(x_{s},u_{s})$ is the one that minimizes the performance measure $J(x,u)$ defined for the population of agents.
+Assumption 3 (Local equilibrium cost bound).
+There exists a constant $\rho>0$ and $\alpha_{\rho}\in\mathcal{K}$ such that $\forall(x,u)\in\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})\cap\mathcal{Z}_{N}(x_{s})$ and $\forall k\in\mathbb{N}$ it holds:
+$\ell(x_{s},u_{s})\leq\ell(x_{k},u_{k})-\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ (22)
+where $\mathcal{Z}_{N}(x_{s})$ is the compact set of all feasible (game) trajectories of length $N$ initialized at $x_{s}$ as in (4).
+In principle, one could also consider $\mathcal{S}^{\rho_{1}\text{-\tiny GNE}}_{N}(\mathbf{x})$ with a smaller “suboptimality” gap $\rho_{1}\leq\rho$ which still ensures that Assumption 3 holds for all trajectories in $\mathcal{S}^{\rho_{1}\text{-\tiny GNE}}_{N}(\mathbf{x})$.
 Assumption 3 can be verified by employing the game-theoretic counterpart of Fiacco-like sensitivity analysis.
-Specifically, in the GNE setting formulated as quasi-variational inequalities, local uniqueness results exist under different sets of assumptions, see [Alphonse et al., 2021, Ch.
-3] and [Noor et al., 2025, Lem.
+Specifically, in the GNE setting formulated as quasi-variational inequalities, local uniqueness results exist under different sets of assumptions, see [1, Ch.
+3] and [49, Lem.
 7].
-For variational GNEs, sensitivity analysis of variational inequalities can be exploited [Facchinei and Pang, 2009, Prop.
+For variational GNEs, sensitivity of variational inequalities can be exploited [25, Prop.
 12.14].
 The following result confirms a converse relation between turnpike and strict dissipativity.
-Theorem 4 (Turnpike $\Rightarrow$ str. dissipativity).
-Suppose at $(x_{s},u_{s})$ Assumption 1 and the turnpike property as per Definition 4 holds for the GNEP (2), then the GNEP is strictly dissipative for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ with respect to $(x_{s},u_{s})$.
-Proof 3.3.
-Consider $\alpha_{\rho}\in\mathcal{K}$ from Assumption 1 and note that for any $\rho$ there exist an $N\in\mathbb{N}$ such that $Q_{\varepsilon}>0$ as defined in (14).
-Consequently, we choose $\rho\geq\varepsilon$ and thus whenever we are in $Q_{\varepsilon}$ we can exploit Assumption 1.
-Recall the available storage from Theorem 1
-$\tilde{\Lambda}_{a}(\mathbf{x})=\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$ (18)
+Theorem 4 (Turnpike $\Rightarrow$ strict dissipativity).
+Suppose that the stage cost $\ell$ is continuous, that at $(x_{s},u_{s})$ Assumption 3, and the turnpike property as per Definition 4 holds for the GNEP (2).
+Then, there exists some $\alpha\in\mathcal{K}$, such that the GNEP is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ with respect to $(x_{s},u_{s})$.
+Moreover, there exists a storage function $\Lambda$ bounded on $\mathbb{X}_{\infty}(\mathbb{X}_{0})$.
+Proof 3.4.
+Consider $\rho$ for which Assumption 3 holds and the corresponding $\alpha_{\rho}\in\mathcal{K}$.
+The turnpike property in (19) states that given $\rho$ [^1] To improve readability of the proof, we choose $\varepsilon=\rho$ but one may also consider a smaller region around $x_{s}$ (i.e., $\varepsilon\leq\rho$).
+there exists a horizon length $N\in\mathbb{N}$ for which the GNE trajectory spends nonzero discrete time steps $Q_{\rho}>0$ within $\mathcal{B}_{\varepsilon}(x_{s})$.
+Further, recall the definition of the available storage from Theorem 1
+${\Lambda}_{\alpha_{\rho}}(\mathbf{x})=\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$ (23)
 evaluated along the game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
-Temporarily, consider some $N\in\mathbb{N}$ and a feasible initial condition $\mathbf{x}$.
-We split the time horizon into $Q_{\varepsilon}(N,\mathbf{x})$, the trajectory points close to $(x_{s},u_{s})$ and $Q_{out}(N,\mathbf{x})=\{0,\dots,N-1\}\backslash Q_{\varepsilon}(N,\mathbf{x})$ points outside of ball of radius $\varepsilon$.
-Evaluating (18) on $Q_{out}(N,\mathbf{x})$ gives
+Here, without loss of generality, we use the $\mathcal{K}$-class function $\alpha_{\rho}$ from Assumption 3.
+Temporarily, consider some $N\in\mathbb{N}$ and a feasible initial condition $\mathbf{x}$. We split the time horizon into $Q_{\rho}(N,\mathbf{x})$, the trajectory points close to $(x_{s},u_{s})$, and
+$Q_{out}(N,\mathbf{x}):=\{0,\dots,N-1\}\backslash Q_{\rho}(N,\mathbf{x}),$
+i.e. the points outside of $\mathcal{B}_{\rho}(x_{s})$. Evaluating (23) on $Q_{out}(N,\mathbf{x})$ gives
 $\sum_{k\in Q_{out}(N,\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$
 […]
-where $\hat{\alpha}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ and similarly $\hat{s}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}s(x_{k},u_{k})$.
-The Turnpike property implies that any trajectory, independent of $N$ and $\mathbf{x}$, only spends a finite amount of time outside of ball the of radius $\varepsilon$ of $x_{s}$ which allows us to establish the following bound:
+where
+$\hat{\alpha}=\sup_{(x,u)\in\mathcal{Z}_{N}(\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$
+and similarly $\hat{s}=\sup_{(x,u)\in\mathcal{Z}_{N}(\mathbf{x})}s(x_{k},u_{k})$.
+The turnpike property implies that any trajectory, independent of $N$ and $\mathbf{x}$, only spends a finite amount of time outside of the ball $\mathcal{B}_{\rho}(x_{s})$.
+This allows us to establish the bound
 […]
-Next, we evaluate the second part of the trajectory for $Q_{\varepsilon}(N,\mathbf{x})$ we have that $\|x_{k}-x_{s}\|\leq\rho,\;\forall k\in Q_{\varepsilon}(N,\mathbf{x})$ and given Assumption 1 it holds that:
-$\sum_{Q_{\varepsilon}(N,\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})\leq 0$ (21)
-Combining (3.3) and (21) we get that
+where due to (19) $\bar{Q}_{out}\geq\frac{C}{\alpha_{\rho}(\rho)}$ is some constant independent of $N$ and $\mathbf{x}$.
+Next, we evaluate the second part of the trajectory.
+For $Q_{\rho}(N,\mathbf{x})$ we have that due to the turnpike property $\forall k\in Q_{\rho}(N,\mathbf{x})$ it holds that $\|x_{k}-x_{s}\|\leq\rho$.
+Assumption 3 holds $\forall k\in Q_{\rho}(N,\mathbf{x})$ by choice of $\rho$.
+Thus we can construct the following inequality:
+$\sum_{k\in Q_{\rho}(N,\mathbf{x})}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})\leq 0$ (26)
+where we use the fact that $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ . Combining (3.4) and (26) we get that
+$\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$
 […]
 […]
-3.3 Optimal operation at steady-state GNE
+Lastly, the existence of a bounded storage function follows from the observation that the available storage is bounded on $\mathbb{X}_{\infty}(\mathbb{X}_{0})$, cf. Appendix A.
+This theorem highlights the close relation of dissipativity and turnpike properties in GNEPs.
+Theorems 3 and 4 show that under rather mild conditions, strict GNEP dissipativity in the sense of Definition 3 and the turnpike property from Definition 4 are equivalent.
+Corollary 5 (Turnpike $\Leftrightarrow$ strict dissipativity).
+Consider the GNEP (2) and let Assumptions 13 hold.
+Suppose that the stage cost $\ell$ is continuous and let $\mathbb{X}_{N}(\mathbb{X}_{0})=\mathbb{X}_{0}$.
+Then the following two statements are equivalent:
+  1. (i)
+    The GNEP (2) is strictly dissipative with bounded storage (Definition 3).
+  2. (ii)
+    The GNEP (2) has the measure turnpike property (Definition 4).
+Due to Theorems 3 and 4, the proof is straightforward and thus omitted.
+The condition $\mathbb{X}_{N}(\mathbb{X}_{0})=\mathbb{X}_{0}$ is imposed to match the set of initial conditions for which Assumption 1 holds with the domain of the available storage.
+This condition is satisfied if from any point on a GNEP solution trajectory the turnpike $(x_{s},u_{s})$ can be reached in finitely many steps.
+3.2 Optimal game operation at steady-state GNEs
 We have established in the previous section under which conditions the open-loop trajectory will spend most of its time at the turnpike, the steady-state GNE.
-However, this raises the question if the pair $(x_{s},u_{s})$ is also optimal in terms of the entire population performance measure $J(x,u)$.
-To investigate this, we define optimal steady state operation and suboptimal operation, following [Müller et al., 2015, Def.
+However, this raises the question if the pair $(x_{s},u_{s})$ is also the best performing GNE (of potentially very many GNEs) in terms of the entire population performance measure $J_{N}(x,u)$ from (13).
+To investigate this, we define optimal steady state operation and suboptimal operation, following [47, Def.
 2].
-Definition 6 (Optimal operation of RHG).
-The GNEP (2) is optimally operated at steady-state, if for each $\mathbf{x}\in\mathbb{X}_{0}$ and each game pair $(x,u)\in\mathcal{S}^{\text{\tiny NE}}_{N}(\mathbf{x})$ the following holds $\forall k\in\mathbb{I}_{\geq 0}$:
-$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}\geq\ell(x_{s},u_{s}).$ (22)
-it is suboptimally operated off steady state if (22) holds in a strict sense.
+Definition 6 (Optimal game operation).
+The GNEP (2) is optimally operated at a steady-state GNE, if for each $\mathbf{x}\in\mathbb{X}_{0}$ and each game pair $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ the following holds $\forall k\in\mathbb{I}_{\geq 0}$
+$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}\geq\ell(x_{s},u_{s})$ (28)
+for $(x_{s},u_{s})\in\mathcal{S}_{s}^{\text{\tiny GNE}}(\mathbf{x})$. It is suboptimally operated off steady state GNE if (28) holds in a strict sense.
 Optimal steady-state GNE operation implies that no feasible game pair of (2) can have a better (asymptotic) average performance than the performance of the steady-state GNE $(x_{s},u_{s})$, and suboptimal operation off steady-state means the performance is strictly worse or “passes by” $x_{s}$ infinitely often.
@@
 Proposition 1.
-Suppose the set $\mathcal{S}_{\infty}^{GNE}(\mathbf{x})\not$$\emptyset$, then if GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ following Definition 3, it is optimally operated at the steady-state GNE $(x_{s},u_{s})$ and suboptimally operated off steady state.
-Proof 3.4.
-The proof follows along the lines of [Prop. 6.4][Angeli et al., 2012]. First, we note that $\sup_{x\in\mathbb{X}_{0}}|\Lambda(x)|<\infty$ which allows to state the following equality
+Suppose that the set $\mathcal{S}_{\infty}^{GNE}(\mathbf{x})\not$$\emptyset$, then if GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ following Definition 3 and with bounded storage, it is optimally operated at the steady-state GNE $(x_{s},u_{s})$ and suboptimally operated off steady state.
+Proof 3.5.
+The proof follows along the lines of [Prop.
+6.4][3].
+First, we note that $\sup_{x\in\mathbb{X}_{0}}|\Lambda(x)|<\infty$ giving a bound on the storage function and allowing us to state the following equality
 $0$ $=\lim_{N\to+\infty}\frac{\Lambda(x_{N})-\Lambda(x_{0})}{N}$
@@
 $\leq\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)}{N}-\ell(x_{s},u_{s})$
-From which it follows that
+Note that switching to $\liminf$ is necessary as the sequence $\ell(x_{k},u_{k})$ may not be converging.
+Using superadditivity [^2] Superadditivity: $\liminf\limits_{n\to\infty}(a_{n}+b_{n})\geq\liminf\limits_{n\to\infty}a_{n}+\liminf\limits_{n\to\infty}b_{n}$ of the $\liminf$ and noting that $\ell(x_{s},u_{s})$ is constant we obtain
 $\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}$
@@
 proving optimal operation at steady state.
-4 Optimality Interpretation of GNEs
-In this section we aim to draw a connection between the GNE solution of (2) and the KKT conditions of optimal control problems.
-In order to do so, we state the following assumptions on the functions in the GNEP (2).
+To the best of our knowledge, Proposition 1 is the first characterization of GNEs from an optimality perspective.
+It shows that if the group of agents converges to $(x_{s},u_{s})$ (i.e., strict dissipativity holds), then this point is the best possible (hence optimal) steady-state GNE for the population in terms of cost.
+[^3] Note that such a characterization exists for potential games [46] that have a direct connection to optimization problems through the potential function.
+Yet, in (2) we consider a general class of GNEPs.
+Figure 2: Overview of implications between strict dissipativity, turnpike, and optimal operation at the steady-state GNE.
+4 Value Function and KKT Interpretation
+In this section we aim to draw a connection between the GNE solution of (2) and their KKT conditions. In order to do so, we state the following assumptions on the functions in the GNEP (2).
 Assumption 4 (Differentiable problem data).
-For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ and constraint function $g(\cdot,\cdot,u^{-v})$ and the dynamics $f(\cdot,\cdot,u^{-v})$ are continuously differentiable in $x$ and $u^{v}$.
+For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J_{N}^{v}(\cdot,\cdot,u^{-v})$, the constraint functions $g(\cdot,\cdot,u^{-v}),h^{v}(\cdot),$ and the dynamics $f(\cdot,\cdot,u^{-v})$ are continuously differentiable in $x$ and $u^{v}$.
 Assumption 5 (Convexity).
-For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ is convex and the set $\mathcal{Z}^{v}(\cdot,u^{-v})$ is closed and convex.
+For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ is convex and the set $\mathcal{Z}_{N}^{v}(\cdot,u^{-v})$ is closed and convex.
 The convexity assumption is required to ensure that points which solve the individual agent’s KKT system are not merely stationary points (local minima, saddle points).
-KKT conditions of GNEs and relevant constraint qualifications have been studied in detail in [Bueno et al., 2019].
+KKT conditions of GNEs and relevant constraint qualifications have been studied in detail in [9].
 At each time step $k=0,\ldots,N-1$ consider
-$L^{v}_{k}=\begin{bmatrix}1\\ \lambda^{v}_{k+1}\\ \mu_{k}^{v}\\ \nu_{k}^{v}\end{bmatrix}^{\top}\begin{bmatrix}\ell^{v}(x_{k},u_{k},u^{-v}_{k})\\ f(x_{k},u_{k}^{v},u_{k}^{-v})-x_{k+1}\\ g(x_{k},u_{k}^{v},u^{-v}_{k})\\ G^{v}u_{k}^{v}-h^{v}\end{bmatrix}$.
+$L^{v}_{k}=\begin{bmatrix}1\\ \lambda^{v}_{k+1}\\ \mu_{k}^{v}\\ \nu_{k}^{v}\end{bmatrix}^{\top}\begin{bmatrix}\ell^{v}(x_{k},u_{k},u^{-v}_{k})\\ f(x_{k},u_{k}^{v},u_{k}^{-v})-x_{k+1}\\ g(x_{k},u_{k}^{v},u^{-v}_{k})\\ h^{v}(u_{k}^{v})\end{bmatrix}$.
 The Lagrangian of each agent associated with (2) can then be written as
@@
 $\lambda^{v}_{k}$ $=\ell_{x}^{v}+g_{x}^{\top}\mu^{v}_{k}+f_{x}^{\top}\lambda_{k+1}^{v},\quad$ (30b)
-$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\mu^{v}_{k}+f_{u^{v}}^{\top}\lambda_{k+1}^{v}+(G^{v})^{\top}\eta_{k}^{v},$ (24c)
+$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\mu^{v}_{k}+f_{u^{v}}^{\top}\lambda_{k+1}^{v}+h_{u^{v}}^{\top}\eta_{k}^{v},$ (30c)
 subject to the initial and boundary conditions
@@
 $0\leq\mu_{k}^{v}$ $\quad\perp\quad-g^{v}(x_{k},u_{k}^{v},u_{k}^{-v})\geq 0,$ (30e)
-$0\leq\eta_{k}^{v}$ $\quad\perp\quad h^{v}-G^{v}u^{v}_{k}\geq 0.$ (24f)
-A well-established result from the GNE literature connects this KKT system to GNEs. The following result appeared as Thm. 4.6 in [Facchinei and Kanzow, 2009].
+$0\leq\eta_{k}^{v}$ $\quad\perp\quad-h^{v}(u^{v}_{k})\geq 0.$ (30f)
+A well-established result from the GNE literature connects this KKT system to GNEs, cf. [24, Thm. 4.6].
 Theorem 6 (KKT conditions for GNEPs).
 Given Assumption 4 then the following statements hold for the GNEP (2):
-  1. (a)
+  1. (i)
     Let $(x^{*},u^{*})$ be an equilibrium of the GNEP at which all agents’ subproblems satisfy Slater’s constraint qualification.
 Then, $\forall v\in\mathcal{V}:\exists\,(\lambda^{v*},\mu^{v*},\eta^{v*})$ that together with $(x^{*},u^{*})$ solve the KKT system (30).
-  2. (b)
-    If $(x^{*},u^{*})$ and $(\lambda^{v*},\mu^{v*},\eta^{v*})_{v\in\mathcal{V}}$ solve the KKT system (24) and Ass. 3 holds, then $(x^{*},u^{*})$ is an equilibrium point of the GNEP.
-Similar to [Faulwasser and Zanon, 2018], we study when a primal-dual solution to the steady-state GNEP (10) is also a solution for the dynamic counterpart (2).
+  2. (ii)
+    If $(x^{*},u^{*})$ and $(\lambda^{v*},\mu^{v*},\eta^{v*})_{v\in\mathcal{V}}$ solve the KKT system (30) and Ass. 5 holds, then $(x^{*},u^{*})$ is an equilibrium point of the GNEP.
+Similar to [29], we study when a primal-dual solution to the steady-state GNEP (11) is also a solution for the dynamic counterpart (2).
 To this end, we consider the Lagrangian for each agent of the steady-state GNE problem in (11) given by
 $\mathcal{L}_{s}^{v}(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})=\ell^{v}(\bar{x},\bar{u}^{v},\bar{u}^{-v})+(\bar{\mu}^{v})^{\top}\,g(\bar{x},\bar{u}^{v},\bar{u}^{-v})$
-$+(\bar{\lambda}^{v})^{\top}[f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}]+(\bar{\eta}^{v})^{\top}(G^{v}\bar{u}^{v}-h^{v}).$
+$+(\bar{\lambda}^{v})^{\top}[f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}]+(\bar{\eta}^{v})^{\top}h^{v}(\bar{u}^{v}).$
 Computing the usual stationarity condition with respect to $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ gives the following set of conditions
@@
 $\bar{\lambda}^{v}$ $=\ell_{x}^{v}+g_{x}^{\top}\bar{\mu}^{v}+f_{x}^{\top}\bar{\lambda}^{v},$ (31b)
-$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\bar{\mu}^{v}+f_{u^{v}}^{\top}\bar{\lambda}^{v}+(G^{v})^{\top}\bar{\eta}^{v},$ (25c)
+$0$ $=\ell_{u^{v}}^{v}+g_{u^{v}}^{\top}\bar{\mu}^{v}+f_{u^{v}}^{\top}\bar{\lambda}^{v}+h_{u^{v}}^{\top}\bar{\eta}^{v},$ (31c)
 which together with
 $0\leq\bar{\mu}^{v}$ $\quad\perp\quad-g^{v}(\bar{x},\bar{u}^{v},\bar{u}^{-v})\geq 0,$ (31d)
-$0\leq\bar{\eta}^{v}$ $\quad\perp\quad h^{v}-G^{v}\bar{u}^{v}\geq 0,$ (25e)
+$0\leq\bar{\eta}^{v}$ $\quad\perp\quad-h^{v}(\bar{u}^{v})\geq 0,$ (31e)
 form the KKT conditions for (11).
@@
 In the analysis of optimal control problems, the value function plays a fundamental role.
-It allows to conclusively characterize optimality in Dynamic Programming, it can also be linked to the Maximum Principle.
-Specifically, in first-order optimality conditions of the Maximum Principle or its simpler Euler-Lagrange variants the gradient of the value function is related to the co-state/adjoint, i.e., to the dual variable linked to the equality constraints imposed by the dynamics.
-Given that (24) are from a per-agent perspective just first-order (KKT) conditions—though all agents consider the same dynamics and they have to agree on the same sate trajectory—one may wonder what is a GNEP-specific counterpart of the optimal control value function?
+It allows to characterize optimality in Dynamic Programming [7], it can also be linked to the Maximum Principle.
+Specifically, in first-order optimality conditions of the Maximum Principle the gradient of the OCP value function is related to the co-state/adjoint, i.e., to the dual variable linked to the equality constraints imposed by the dynamics.
+Given that (30) are from a per-agent perspective just first-order (KKT) conditions—though all agents consider the same dynamics and state trajectory—one may wonder is there a GNEP-specific counterpart of the optimal control value function?
 To this end, recall the construction of $\ell$ as the sum over all agent-specific costs $\ell^{v}$, cf.
 (12).
-Summing $\ell$ from $k=0$ to k=N1k=N-1 leads to the game value function which measures the performance of the agent population at $\mathbf{x}$
-$\hskip-4.2679ptV_{N}(\mathbf{x}):=\sum_{k=0}^{N-1}\ell(x_{k}^{*},u^{*}_{k})=\sum_{v\in\mathcal{V}}\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$ (26)
-Let $\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}$ denote an open ball of radius $\varepsilon$, centered at $\mathbf{x}$.
+Summing $\ell$ from $k=0$ to k=N1k=N-1 leads to the game value function $V_{N}^{*}:\mathbb{X}_{0}\to\mathbb{R}$ which measures the performance of the agent population at $\mathbf{x}$
+$\hskip-4.2679ptV^{*}_{N}(\mathbf{x}):=\sum_{k=0}^{N-1}\ell(x_{k}^{*},u^{*}_{k})=\sum_{v\in\mathcal{V}}\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$ (32)
 Theorem 7 (Sensitivity characterization).
@@
 $x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$
-we have $\mathcal{S}^{GNE}_{N}(\mathbf{x}+\xi)\not$$\emptyset$, i.e., the GNEP (2) admits a solution such that $V_{N}(\mathbf{x})$ is locally differentiable at $\mathbf{x}$.
+we have $\mathcal{S}^{GNE}_{N}(\mathbf{x}+\xi)\not$$\emptyset$, i.e., the GNEP (2) admits a solution such that $V^{*}_{N}(\mathbf{x})$ is locally differentiable at $\mathbf{x}$.
 Then
-$\nabla V_{N}(\mathbf{x})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0},$ (27)
+$\nabla V^{*}_{N}(\mathbf{x})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0},$ (33)
 where $\lambda^{v}_{0}$ corresponds to the KKT solution tuple $(x,u,\lambda^{v},0,0)$ with $(x,u)\in\mathcal{S}^{GNE}_{N}(\mathbf{x})$.
 Proof 4.1.
-Consider the GNEP (24) without inequality constraints and subject to the perturbed initial condition
-$x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}.$
-For all $v\in\mathcal{V}$, let $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ denote the solution to the KKT conditions (24) with perturbed initial condition, which for all $\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$ are optimal due to Theorem 5.
-Clearly, it holds that the solution for the unperturbed initial condition is given by
+Consider the GNEP (30) without inequality constraints and subject to the perturbed initial condition $x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}$.
+For all $v\in\mathcal{V}$, let $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ denote the solution to the KKT conditions (30) with perturbed initial condition, which for all $\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$ are dynamic GNEs due to Theorem 6.
+The solution for the unperturbed initial condition is
 $(x^{*},u^{v,*},\lambda^{v})=(x(0),u^{v}(0),\lambda^{v}(0)).$
 Computing the sensitivity of the Lagrangian (29) with respect to $\xi$ at $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ gives
-$\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}\left(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)\right)=$ $\phantom{+}\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial x_{k}}\dfrac{\partial x_{k}}{\partial\xi}$
-$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial u^{v}_{k}}\dfrac{\partial u^{v}_{k}}{\partial\xi}$
-$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial\lambda^{v}_{k}}\dfrac{\partial\lambda^{v}_{k}}{\partial\xi}$
-$+\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi},$
+$\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}\left(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)\right)=\\ \sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial x_{k}}\dfrac{\partial x_{k}}{\partial\xi}+\dfrac{\partial L_{k}^{v}}{\partial u^{v}_{k}}\dfrac{\partial u^{v}_{k}}{\partial\xi}+\dfrac{\partial L_{k}^{v}}{\partial\lambda^{v}_{k}}\dfrac{\partial\lambda^{v}_{k}}{\partial\xi}+\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi},$
 where $\frac{\partial x_{k}}{\partial\xi}$ is the sensitivity of state $x_{k}(\xi)$ to the perturbation of the initial condition $x_{0}=\mathbf{x}+\xi$.
@@
 Hence, for all $v\in\mathcal{V}$, we obtain
-$\left.\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))\right|_{\xi=0}$ $=\left.\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi}\right|_{\xi=0}$
+$\hskip 0.0pt\left.\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))\right|_{\xi=0}$ $=\left.\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi}\right|_{\xi=0}$
 $=\lambda^{v}(0)^{\top}.$ (34)
-Observe that the definition of $V_{N}(\mathbf{x})$ in (26) implies
-$V_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi).$
+Observe that the definition of $V^{*}_{N}(\mathbf{x})$ in (32) implies
+$V^{*}_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
 differentiating on both sides with respect to $\xi$ gives
-$\dfrac{\mathrm{d}}{\mathrm{d}\xi}V_{N}(\mathbf{x}+\xi)=\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
-Evaluation at $\xi=0$ and using (4.1) proves the assertion.
+$\dfrac{\mathrm{d}}{\mathrm{d}\xi}V^{*}_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
+Setting $\xi=0$ and using (4.1) proves the assertion.
+Sensitivity characterization of the value function is an important step towards a more general characterization of value functions in GNEPs with self-interested agents and allows to draw connections to the storage function.
 Figure 3: Finite-horizon and steady-state GNEP with GNE-KKT system
 4.2 Sensitivity analysis of the storage function
-The next result translates a crucial observation made in the context of turnpike analysis of optimal control problems to GNEPs.
-In some sense, the next result provides a storage function counterpart to Theorem 6.
+The next result translates a crucial observation made in the context of turnpike analysis of OCPs to GNEPs providing a storage function counterpart to Theorem 7.
 Theorem 8 (Storage gradient).
@@
   1. (i)
-    Any primal-dual solution $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ of (25) constitutes a steady-state solution of (24a24c).
+    Any primal-dual solution $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ of (31) constitutes a steady-state solution of (30a30c).
   2. (ii)
-    Let $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v}=0,\bar{\nu}^{v})$ solve (25) such that $(\bar{x},\bar{u})\in\Pi\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ for some $\mathbf{x}\in\mathcal{X}_{0}$.
+    Let $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v}=0,\bar{\nu}^{v})$ solve (31) and $\bar{x}\in\mathbb{X}_{N}(\mathbb{X}_{0})$ holds.
 Then, for any storage function $\Lambda$ differentiable at $\bar{x}$, it holds that
@@
 $0=\ell_{x}-\left(f_{x}^{\top}-I\right)\nabla\Lambda(\bar{x})$ (36)
-with $\ell$ from (11).
-Next, we sum (25b) over $v\in\mathcal{V}$ and obtain
+with $\ell$ from (12). Now sum (31b) over $v\in\mathcal{V}$ to obtain
 $0=\sum_{v\in\mathcal{V}}\ell_{x}^{v}+\left(f_{x}^{\top}-I\right)\lambda^{v}=\ell_{x}+\left(f_{x}^{\top}-I\right)\sum_{v\in\mathcal{V}}\bar{\lambda}^{v}$
 where we use (12) and the fact that $\bar{\mu}^{v}=0$. Comparison with (36) directly gives the assertion.
-Note that Theorem 7 (i) states that (25a25c) is a fixed point for (24a24c) but not for (2) as this would also require the boundary conditions (24d) to hold.
+Note that Theorem 8 (i) states that (31a31c) is a fixed point for (30a30c) but not for (2) as this would also require the boundary conditions (30d) to hold.
 Moreover, the negative sign in (35) stems from the steady-state constraint $f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}=0$ in (11).
-If this constraint is equivalently written as $\bar{x}-f(\bar{x},\bar{u}^{v},\bar{u}^{-v})=0$, then corresponding Lagrange multiplier flips sign and the minus in (29) can be dropped.
+If this constraint is equivalently written as $\bar{x}-f(\bar{x},\bar{u}^{v},\bar{u}^{-v})=0$, then the corresponding Lagrange multiplier flips sign and the minus in (35) can be dropped.
 If $\bar{\mu}^{v}\not=0$, there exist active per-agent constraints in the steady-state problem (11). The previous result can then be extended to
@@
 where $(\cdot)^{\dagger}$ denotes the Moore-Penrose inverse.
-The above result shows two important links between the GNEP (2) and its steady-state counterpart (10): The optimality systems of both problems are closely related which resembles insights from optimal control theory [Trélat and Zuazua, 2015, Zanon and Faulwasser, 2018, Faulwasser and Grüne, 2022].
-Moreover, there exists a strong link between the dual variables $\lambda^{v}$ and the storage function $\Lambda$.
-In the context of optimal control problems similar results have been shown in [Damm et al., 2014, Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018].
-Yet, in the context of GNEPs this link is novel, as the dissipativity-based analysis of GNEPs is still in its infancy.
+The above result shows two important links between the GNEP (2) and its steady-state counterpart (11): The optimality systems of both problems are closely related which resembles insights from optimal control theory [54, 60, 28].
+Moreover, there exists a strong link between the dual variables $\lambda^{v}$ and the storage function $\Lambda$ which has been previously shown for OCPs [18, 60, 29].
+Yet, to the best of our knowledge in the context of GNEPs this link is novel.
 We conclude this section with a corollary that evaluates (35) at the turnpike state $x_{s}$.
@@
 where, for all $v\in\mathcal{V}$, $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ is a KKT point solving (31).
-If the GENP exhibits a turnpike at $(x_{s},u_{s})$ and horizon $N$ is sufficiently large, we arrive at a crucial link between the value function (26) and any differentiable storage function
-$\nabla V_{N}(x_{s})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0}\approx\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s})$
-or, for all $v\in\mathcal{V}$, $\lambda^{v}_{0}\approx\lambda_{s}^{v}$.
+If the GNEP exhibits a turnpike at $(x_{s},u_{s})$ and the horizon $N$ is sufficiently large, we arrive at a crucial link between the value function (32) and any differentiable storage function
+$\nabla V^{*}_{N}(x_{s})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0}\approx\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s})$
+or, for all $v\in\mathcal{V}$, $\lambda^{v}_{0}\approx\lambda_{s}^{v}$. The horizon being sufficiently large ensures that the GNEP trajectory converges sufficiently close to $x_{s}$.
 4.3 Turnpike leaving arc and linear end penalties
-A well-known aspect of the turnpike property is the characteristic leaving arc, where state and input trajectories diverge from the turnpike in the final part of the considered time horizon, cf.
+A well-known aspect of the turnpike property is the characteristic leaving arc, where state and input trajectories diverge from the turnpike in the final steps of the horizon, cf.
 Figure 1.
-While the formal Definition 4 does not require that a turnpike solution has a leaving arc, the early works [Dorfman et al., 1958, McKenzie, 1976] and many later results [Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] focus on it.
+While the Definition 4 does not require a leaving arc, the early works [20, 45] and later results [2, 54] focus on it.
 In the context of optimal control, the presence of a leaving arc arises when the values of the states $x$ and the adjoints $\lambda$ close to the turnpike differs substantially from their values at the end of the horizon.
-In optimal control there are three mechanism which generate leaving arcs [Faulwasser and Grüne, 2022]:
+In optimal control there are three mechanisms which generate leaving arcs [28]:
   • Terminal constraints on the state variable such that $x_{N}=x^{N}\not=x_{s}$. Naturally this can be generalized to $x_{N}\in\mathbb{X}_{N}$.
-  • Terminal stage costs $V_{\mathrm{f}}(x)$ which (without any active constraints at $k=N$) imply that the transversality condition $\lambda_{N}=\nabla V_{\mathrm{f}}(x)$.
-  • Or the presence of state constraints which render it cheap to let a system drift towards the constraint boundary at the end of the horizon.
-Interestingly, the same leaving arc phenomenon can be observed in GNEPs, see, e.g., the supply chain example [Hall et al., 2024a].
+  • Terminal stage costs $V_{\mathrm{f}}(x)$ which (without any active constraints at $k=N$) imply the transversality condition $\lambda_{N}=\nabla V_{\mathrm{f}}(x)$.
+  • Or the presence of state constraints and a cost which renders it cheap to let a system drift towards the constraint boundary at the end of the horizon.
+Interestingly, the same leaving arc phenomenon can be observed in GNEPs, see, e.g., the supply chain example [38].
 Moreover, the optimality interpretation of GNEPs—and in particular the structure of the KKT conditions, which is very similar to classic optimal control—suggest that the same mechanisms generate a leaving arc.
-Inspired by the OCP analysis in [Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018], the next result shows how to suppress leaving arcs.
+We have established in Proposition 1 that under a strict dissipativity assumption the steady-state GNE in (11).
+Thus, it is desirable to enforce that agents not only converge to $(x_{s},u_{s})$ but remain there.
+Inspired by the OCP analysis in [29], the next result shows how to suppress leaving arcs in GNE trajectories.
 Proposition 10 (GNEPs without leaving arcs).
-Consider the GNEP (2) and let Assumptions 23 hold. Let $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ be a KKT point solving (25). Consider the following two statements:
+Consider the GNEP (2) and let Assumptions 45 hold. Let $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ be a KKT point solving (31). Let one of the following two statements hold:
   1. (i)
@@
     $\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})+V_{\mathrm{f}}^{v}(x_{N})$
-    and
-    $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}.$
+    and $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$.
   2. (ii)
@@
 Proof 4.3.
-As per Theorem 7, part (i,) the steady-state primal-dual tuple $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ from (25) is a steady state of (24a24c).
-Moreover, using $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ to construct an $N$-step constant trajectory also the complementary slackness and the dual feasibility conditions (24e-24f) are satisfied as they correspond point-wise in time to their counterparts in (25).
-It remains to analyze (24d). For case (i), notice that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ in (23) has an additional term
-$V_{\mathrm{f}}^{v}(x_{N})$
-which, by standard KKT analysis leads to the modified boundary condition
-$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda_{N}=\nabla V_{\mathrm{f}}^{v}(x_{N})+g_{x}^{\top}\mu_{N}^{v}.$ (24d-i)
+As per Theorem 8, part (i) the steady-state primal-dual tuple $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ from (31) is a steady state of (30a30c).
+Moreover, using $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ to construct an $N$-step constant trajectory also the complementary slackness and the dual feasibility conditions (30e-30f) are satisfied as they correspond point-wise in time to their counterparts in (31).
+It remains to analyze (30d).
+For case (i), notice that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ in (29) has an additional term $V_{\mathrm{f}}^{v}(x_{N})$ which using standard KKT analysis leads to the modified boundary condition
+$x_{0}=\mathbf{x}\quad\text{and}\quad\lambda_{N}=\nabla V_{\mathrm{f}}^{v}(x_{N})+g_{x}^{\top}\mu_{N}^{v}.$ (30d-i)
 Recall that the assertion sets $\mathbf{x}=x_{s}$.
-From the condition that $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$ it then immediately follows that the $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies $x_{0}=x_{s}$ and (24d-i).
-For case (ii), observe that the Lagrangian $\mathcal{L}^{v}(x,u,\lambda^{v},\mu^{v},\nu^{v})$ now includes the additional term
-$\sigma^{\top}(x_{N}-x_{s}),$
-where $\sigma\in\mathbb{R}^{n}$ is the additional multiplier for the point-wise terminal constraint. Once more, standard KKT analysis leads to the modified boundary condition
-$x_{0}=\mathbf{x}\quad\text{and}\quad x_{N}=x_{s}.$ (24d-ii)
+From the condition that $\nabla V_{\mathrm{f}}^{v}(x_{s})+g_{x}^{\top}\mu_{N}^{v}=\lambda_{s}^{v}$ it then immediately follows that the $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies $x_{0}=x_{s}$ and (30d-i).
+For case (ii), observe that the Lagrangian $\mathcal{L}^{v}$now includes the additional term $\sigma^{\top}(x_{N}-x_{s})$, where $\sigma\in\mathbb{R}^{n}$ is the additional multiplier for the point-wise terminal constraint.
+Once more, standard KKT analysis leads to the modified boundary condition
+$x_{0}=\mathbf{x}\quad\text{and}\quad x_{N}=x_{s}.$ (30d-ii)
 The $N$-step constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},$ $\mu_{s}^{v},\nu_{s}^{v})$ satisfies these boundary conditions as $\mathbf{x}=x_{s}$.
-The preceding analysis has shown that the constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v})$ satisfies the KKT conditions (24).
+The preceding analysis has shown that the constant trajectory staying at $(x_{s},u_{s}^{v},\lambda_{s}^{v},\mu_{s}^{v},\nu_{s}^{v})$ satisfies the KKT conditions (30).
 Theorem 6 then directly gives the assertion.
 This finishes the proof.
-A first connection between the KKT system and the role of the terminal constraint on the dual variable was drawn in [Hall et al., 2024b, Prop.
+A first connection between the KKT system and the role of the terminal constraint on the dual variable was drawn in [35, Prop.
 3] and the proof therein.
 Specifically, it was demonstrated that with a terminal constraint $x_{N}=x_{s}$ and additional assumptions on the cost, the steady-state GNEP $(x_{s},u_{s})$ is the unique fixed point of the finite-horizon GNEP-KKT system and the unique equilibrium point of a closed-loop system resulting from a receding-horizon implementation of the GNE.
+Next, we design a linear end penalty which suppresses the leaving arc.
 Corollary 11 (Linear end penalty).
-Let the conditions of Proposition 9 hold and suppose that $\mu_{s}^{v}=0$ for all $v\in\mathcal{V}$. Then
+Let the conditions of Proposition 10 (i) hold and suppose that $\mu_{s}^{v}=0$ for all $v\in\mathcal{V}$. Then
 $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v},\quad\forall v\in\mathcal{V}$
@@
 $(x_{s},u_{s})\equiv:(x,u)=(x^{*},u^{*})\in\mathcal{S}_{N}^{GNE}(x_{s}).$
-Indeed, quite similarly to the optimal control setting [Zanon and Faulwasser, 2018], also for GNEPs (2), the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ admits a primal and a dual interpretation:
+Indeed, quite similarly to the optimal control setting [60], also for GNEPs (2), the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ admits a primal and a dual interpretation:
   • It directly specifies the boundary or terminal constraint on the dual variable $\lambda^{v}_{N}$.
-Interestingly, the point-wise boundary condition on $\lambda^{v}_{N}$ is an optimality conditions but it does not imply any reachability requirement of the actual primal state dynamics $x^{+}=f(x,u)$.
+Interestingly, the point-wise boundary condition on $\lambda^{v}_{N}$ is an optimality condition but it does not imply any reachability requirement of the actual primal state dynamics $x^{+}=f(x,u)$.
   • Via a converse telescopic sum argument $V_{\mathrm{f}}^{v}(x)=\lambda_{s}^{\top}x$ can be pulled into the stage cost
     $\tilde{\ell}^{v}(x,u^{v},u^{-v})=\ell^{v}(x,u^{v},u^{-v})+\lambda_{s}^{\top}\left(f(x,u^{v},u^{-v})-x\right).$
-    This trick is called rotating the stage cost in economic MPC [Diehl et al., 2011, Angeli et al., 2012].
-The linear rotation with $\lambda_{s}$ implies that the steady constraint in (10) is only weakly active.
-This geometric modification of the stage cost with $\lambda_{s}$ has been called gradient correction in economic MPC, cf.
-[Zanon and Faulwasser, 2018].
+    This is called rotating the stage cost in economic MPC [19, 3].
+The linear rotation with $\lambda_{s}$ implies that the steady-state constraint in (11) is only weakly active.
+It is also called gradient correction, cf.
+[60].
 4.4 Learning the linear end penalty
-One may ask about the pros and cons of the different approaches that allow suppressing the leaving arc of turnpike solutions.
-Indeed, the point-wise terminal constraint $x_{N}=x_{s}$ requires solving the steady-state GNEP (10) as does the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$.
-In settings where the cost function changes occasionally—e.g.
-due to rare parametric updates (prices, etc.)—it is desirable to avoid the extra computational effort.
-Also, the primal terminal constraint $x_{N}=x_{s}$ jeopardize feasibility in case of short horizons.
-In the context of suppressing turnpike leaving arcs in OCPs, [Faulwasser and Zanon, 2018, Sec. 4] proposed an adaptive strategy for the terminal penalty.
-Algorithm 1 summarizes how this strategy may be transferred to GNEPs.
+Both the point-wise terminal constraint $x_{N}=x_{s}$ and the linear end penalty approach $V_{\mathrm{f}}^{v}(x)=x^{\top}\lambda_{s}^{v}$ require solving the steady-state GNEP (11) in advance.
+Yet, in settings where the cost function changes occasionally (e.g.
+due to parametric updates such as prices) it would be desirable to avoid the extra computational effort.
+In the context of suppressing turnpike leaving arcs in OCPs, [29, Sec.
+4] proposed an adaptive strategy for the terminal penalty.
+Algorithm 1 summarizes how this strategy can be transferred to GNEPs.
 In Step 2 the GNEP (2) is solved considering the linear end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, where the initialization $p_{0}=0$ is used.
 Then, based on our turnpike insights, for all $v\in\mathcal{V}$, an approximation of the turnpike dual variable $\lambda^{v}_{s}$ is obtained by $p_{i}^{v}=\lambda^{v}(\frac{N}{2})$ in Step 5.
-The GNEP is resolved using the updated end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$.
-The algorithm ends if either the change in $\|p_{i}^{v}-p_{i+1}^{v}\|$ becomes small or an iteration limit is reached.
+The GNEP is recomputed using the updated end penalty $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$.
+The algorithm ends if either the change in $\|p_{i}^{v}-p_{i-1}^{v}\|$ becomes small or an iteration limit is reached.
 Put differently, the algorithm builds upon the observation that whenever the states and inputs are close to their turnpike values also the dual variables stay close to their turnpike values.
-It is worth remarking that each agent can learn its specific linear end penalty.
+It is worth noting that each agent can learn its specific linear end penalty.
 One may even learn this end penalty for some agent while others still exhibit leaving arcs.
-While the formal convergence analysis is beyond the scope of the present paper, the next section will present numerical results demonstrating the efficacy of the proposed approach.
-Algorithm 1 Learning $\lambda_{s}^{v}$ [Faulwasser and Zanon, 2018]
+While the formal convergence analysis of Algorithm 1 is beyond the scope of this work , the next section will present numerical results demonstrating its efficacy.
+Algorithm 1 Learning the linear end penalty $\lambda_{s}^{v}$
 1:GNEP (2) with horizon $N=2\hat{N}$, initial condition $\mathbf{x}$, and $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$, $i=0,p_{0}=0$, $i_{max}\in\mathbb{N}$, $\Delta_{0}=\infty$, $\varepsilon>0$
 […]
-3:while $i\leq i_{max}$ OR $\Delta_{i}\leq\varepsilon$ do
+3:while $i\leq i_{max}$ AND $\Delta_{i}\geq\varepsilon$ do
 4:  Solve GNEP (2) with $V_{\mathrm{f}}^{v}(x)=x^{\top}p^{v}_{i}$
@@
 The reference state is chosen as $x^{\text{ref}}=0.3$.
-Assumptions 2 and 3 are fulfilled by construction.
+Assumptions 4 and 5 are fulfilled by construction.
 The game pairs $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ for different horizon lengths are shown in Figure 4(a) and the corresponding trajectory of dual variables $\lambda^{v}$ in Figure 4(b).
@@
 Since $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and the steady-state GNE in (11) are not necessarily unique, the terminal penalty method requires verifying that $\lambda_{s}$ corresponds to the global turnpike.
-Local turnpike phenomenon in OCPs have been studied in [Krügel et al., 2023], but this is left for future work in the GNE setting.
+Local turnpike phenomenon in OCPs have been studied in [40], but this is left for future work in the GNE setting.
 Figure 5: Open-loop GNE trajectories of (37) with a $\lambda_{s}\,x_{N}$ penalty.
@@
 We extend the close relation between dissipativity and turnpike behaviors known in optimal control to dynamic generalized Nash games.
+The two main ingredients of our analysis are a novel definition of a strict dissipativity property tailored to GNEPs as well as an assumption upper bounding the price of anarchy.
+Moreover, we establish a local variational characterization of the game value function and relate them to the geometry of storage functions.
 Our simulation results show that the linear end penalties designed based on the optimality conditions successfully suppress the leaving arc of the turnpike even in the coupled, non-cooperative, game-theoretic setting.
@@
 Further, we will study closed-loop stability based on turnpike results when GNEs are applied in a receding-horizon fashion as in MPC.
+Appendix A Proof of Theorem 1
+We adapt the original proof of [58] to the setting of dynamic GNEPs and to the strict dissipation inequality (14).
+Henceforth, we use the shorthand $\alpha_{\ell}(\cdot):=\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$.
+Where for simplicity we dropped the $^{*}$ notation.
+[…]
+$\Lambda_{\alpha_{\ell}}(\mathbf{x})\geq\sum_{k=0}^{m}\alpha_{\ell}\left(\cdot\right)-s(x_{k},u_{k})+\Lambda_{\alpha_{\ell}}(x_{m}).$ (38)
+If at $\mathbf{x}=x_{0}$, we have $N^{*}<m$, optimality likewise implies the above inequality. Rearranging (38) directly gives (14).
+[…]
+$\Lambda(\mathbf{x})+\sum_{k=0}^{m}-\alpha_{\ell}\left(\cdot\right)+s(x_{k},u_{k})\geq\Lambda(x_{m})\geq-c>-\infty.$
+The right hand side inequalities follow from the lower boundedness of storage functions $\Lambda$.
+Notice that these bounds hold for all $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
+Hence, we set $m=N$ and rearrange the inequalities such that
+$\Lambda(\mathbf{x})\geq-c+\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x,u)\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}-\sum_{k=0}^{N}-\alpha_{\ell}\left(\cdot\right)+s(x_{k},u_{k}).$
+Here we have used that c-c is not affected by the optimization over the free end time $N$.
+With $x_{0}=\mathbf{x}$, the previous inequality is equivalent to $\Lambda(\mathbf{x})\geq-c+\Lambda_{\alpha_{\ell}}(\mathbf{x})$.
+Since $c\in\mathbb{R}$, $\Lambda_{\alpha_{\ell}}(\mathbf{x})=\infty$ would contradict $\Lambda(\mathbf{x})\in\mathbb{R}$.
+This finishes the proof.∎
+Appendix B Performance bounds from bounded PoA
+Lemma 12.
+Suppose that $0<\nu\leq V^{\diamond}_{N}(\mathbf{x})\leq V<\infty$ is satisfied.
+  • (i)
+    If Assumption 2 holds with $P\in\mathbb{R}$, then
+    $\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})-V^{\diamond}_{N}(\mathbf{x})\leq VP.$ (39)
+  • (ii)
+    If (39) holds with bound VPVP, then Assumption 2 holds with $\frac{V}{\nu}P+1$.
+Proof B.1.
+Part (i):  If Assumption 2 holds, we have
+$\dfrac{1}{V^{\diamond}_{N}(\mathbf{x})}\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq P\leq P+1.$
+Multiplying both sides with $V^{\diamond}_{N}(\mathbf{x})\geq\nu>0$ gives
+$\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq V^{\diamond}_{N}(\mathbf{x})(P+1).$
+Substracting $V^{\diamond}_{N}(\mathbf{x})$ and using $P\cdot V^{\diamond}_{N}(\mathbf{x})\leq PV$ gives the assertion.
+Part (ii):  Consider (39), add $V^{\diamond}_{N}(\mathbf{x})$ on both sides, and divide by $V^{\diamond}_{N}(\mathbf{x})$ to obtain
+$\dfrac{1}{V^{\diamond}_{N}(\mathbf{x})}\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})\leq\dfrac{V}{V^{\diamond}_{N}(\mathbf{x})}P+1$
+The left hand side is bounded from above by $\frac{V}{\nu}P+1$ which finishes the proof.